### v1

# Adversarial thermodynamics

Abstract
In thermodynamics, an agent’s ability to extract work is fundamentally constrained by their environment. Traditional frameworks struggle to capture how strategic decision-making under uncertainty — particularly an agent’s tolerance for risk — determines the trade-off between extractable work and probability of success in finite-scale experiments. Here, we develop a framework for non-equilibrium thermodynamics based on adversarial resource theories, in which work extraction is modelled as an adversarial game for an agent extracting work. Within this perspective, we recast the Szilard engine as a game isomorphic to Kelly gambling, an information-theoretic model of optimal betting under uncertainty — but with a thermodynamic utility function. Extending the framework to finite-size regimes, we apply a risk-reward trade-off to find an interpretation of the Renyi-divergences, in terms of extractable work for a given failure probability. By incorporating risk sensitivity via utility functions, we show that the guaranteed amount of work a rational agent would accept instead of undertaking a risky protocol is given by a Rényi divergence. This provides a unified picture of thermodynamics and gambling, and highlights how generalized free energies emerge from an adversarial setup.
I Introduction
The second law of thermodynamics is a cornerstone of modern physics and has proved successful in describing heat engines, black hole physics, and the thermodynamics of everyday life. Traditionally, for large systems at equilibrium, the second law can be understood in two complementary ways: either as a constraint on average entropy production (or, equivalently, as a decrease in free energy) or as a rule governing which state transitions are physically possible.
In recent decades, advances in nanotechnology and biophysics have made the study of non-equilibrium and small-scale thermodynamics increasingly important. In these regimes—where the thermodynamic limit no longer applies and systems contain only finitely many interacting particles—ensemble averages fail and fluctuations dominate. Deterministic quantities such as free energy must be replaced by probabilistic, protocol-dependent notions, leading to tradeoffs between extractable work and probability of success. The breakdown of determinism has motivated two main approaches to understand the second law of thermodynamics in regimes where standard assumptions break down: stochastic thermodynamics, and the resource-theoretic approach (for a review see [1]).
Stochastic thermodynamics[2, 3] embraces the inherent randomness of small, out-of-equilibrium systems, by treating work as a fluctuating quantity. It describes how entropy on average must increase, while accounting for rare transient decreases. On the other hand, the resource-theoretic approach[4, 5, 6] reformulates the second law as a constraint on possible state transformations. It shifts the focus from averages to operational possibilities, yielding generalized second laws (expressed through Renyi divergences) that quantify how far a state is from thermal equilibrium[7, 8]. While both frameworks are profoundly successful, and their views on the second law can be related[9], neither provides a complete, operational prescription for how an agent’s strategic choices—specifically, their tolerance for risk—directly determine the trade-off between the amount of work they can extract and the probability of successfully obtaining it in a single, finite-scale experiment. In this work, we bridge this gap by treating work extraction from the lens of expected utility theory and decision theory [10].
There have been seminal works forging connections between stochastic thermodynamics and gambling  [11, 12, 13, 14, 15, 16]. One strand of research has implemented specific gambling strategies directly within thermodynamic protocols, as seen in the work extraction experiments of [13, 14]. Concurrently, [12] established formal links between gambling, work extraction, and information flows from an information-thermodynamic perspective. In a separate approach, [16] applied expected utility theory to evaluate thermodynamic processes as lotteries with pre-defined work payoffs. While these works successfully import concepts from one field to the other, our framework provides a deeper analogy between them and reveals a decision-theoretic aspect of thermodynamics. By formally modeling work extraction as an adversarial game implemented by a Szilard-type engine, we demonstrate that generalized free energies emerge from the necessary constraints on rational, strategic play.
We develop a framework based on adversarial resource theories [17], in which an agent must choose a work-extraction protocol while contending with constraints initially imposed on the engine. This perspective casts thermodynamics as a decision-theoretic problem, where the optimal strategy is determined by the agent’s sensitivity to fluctuations. The mathematical structure of this game is formally analogous to the Kelly betting problem from information theory [18], a connection we develop in a companion paper [17]. However, a crucial distinction arises from the physical context: while financial wealth in Kelly gambling grows multiplicatively, thermodynamic work is an additive quantity. This difference dictates distinct classes of rational utility functions—Constant Relative Risk Aversion (CRRA) for gambling versus Constant Absolute Risk Aversion (CARA) for thermodynamics—which in turn shape the optimal strategies and their interpretation.
The paper is structured as follows. In section, II, we formulate a Szilard-type engine as an adversarial setup between an agent (Alice) extracting the work and another agent (Bob) preparing initial constraints, and derive the average extractable work and its connection to the non-equilibrium free energy. In section III, we analyse the risk-reward trade-off through the lens of decision theory, and in particular expected utility theory, where we identify Constant Absolute Risk Aversion (CARA) as the relevant utility class for thermodynamics. This framework allows us to compute both the certainty equivalent (the guaranteed amount of work a risk-sensitive individual would accept instead of undertaking a risky extraction protocol) as Eq (9), and the expected value of the work extracted by a rational individual, Eq (8) for any level of risk aversion. We find that both are parametrized by Renyi divergences. We explain why the former result has no analogue in gambling, while the latter result is related to the gambling result derived in [17]. In Section IV, we approach the problem from an information-theoretic viewpoint on the finite-size regime, where fluctuations dominate. Here, we apply results from concurrent work [19, 17] to show that the work extraction problem reduces to a decision-theoretic problem, introducing a fundamental risk-reward trade-off in which the work extracted if a strategy is successful can be bounded in terms of Renyi divergences $D_{\alpha}$. This gives an operational interpretation to $D_{\alpha}$ for each individual $\alpha$, in terms of the minimal work extraction given some risk tolerance. This extends the result of  [20] who gave an interpretation of the Renyi divergences $D_{0}$ and $D_{\infty}$ in terms of bounds on extracting work in the case of an extremely risk averse or risk seeking agent. In both the expected utility approach and the information-theoretic approach, we give identical and explicit strategies for achieving work extraction given some risk tolerance via Eq (7). Together, all these results provide a unified operational interpretation that bridges stochastic and resource-theoretic views of the second law. Technical details and extended calculations are deferred to the appendices.
II The adversarial Szilard engine
We begin by formalising a thermodynamic work extraction problem as a set-up between adversaries, mirroring the structure of Kelly betting, which we review in appendix A. Consider three players: Bob, who sets initial constraints; Alice, who optimizes work extraction (analogous to a gambler allocating bets), and Charlie, a referee who enforces randomness.
We consider an empty box of volume $V$ which Bob divides into two parts by placing a partition at some position $Q_{B}$. The referee, Charlie, then samples from a binary probability distribution $P_{X}(x)$ and places a molecule on the left or right hand side of the box according to the outcome which Alice and Bob do not know. Alice proceeds to extract work by performing isothermal compressions and expansions on the box, moving the partition to a final position $Q_{A}$ of her choice. The process is illustrated in Fig. 1.
(a) Bob places the partition at some initial position $Q^{B}$
(b) Alice chooses a position $Q^{A}$ to which she moves the partition
(c) When the molecule is on the left hand side of the box, Alice extracts work by moving it to her desired position
(d) When the molecule is on the right hand side of the box, Alice must do work to move the partition to the position of her choice
Figure 1: Schematic of the adversarial Szilard engine.
Without loss of generality, assume that Bob places the partition closer to the leftmost edge of the box, so that the volume on the left hand side can be expressed as $Q_{B}\textit{l}A$, and that Alice moves her partition to a position so that the volume on the left hand side can be expressed as $Q_{A}\textit{l}A$ in terms of fractions $Q_{A}\in(0,1)$ and $Q_{B}\in(0,1)$ of the total length $l$ of the box.
The work extracted by Alice depends on the molecule’s position:
  • When $x=0$ (left-hand side): $w_{0}=k_{B}T\ln(Q^{A}/Q^{B})$
  • When $x=1$ (right-hand side): $w_{1}=k_{B}T\ln\left(\frac{1-Q^{A}}{1-Q^{B}}\right)$
To maintain consistency with the notation of the literature [21], we write $Q^{B}=Q_{X}^{B}(0)$ and $1-Q^{B}=Q_{X}^{B}(1)$, and similarly for Alice. With this notation, the amount of work extracted (which is a random variable) can be written in terms of the positions to which Alice and Bob move the partition (which are not a random variable).
When the process is repeated a large number $n$ of times, the average amount of work $W$ extracted by Alice is given by $W=n\bigg(P_{X}(0)\ln{\bigg(\frac{Q_{X}^{A}(0)}{Q_{X}^{B}(0)}}\bigg)+P_{X}(1)\ln\bigg(\frac{Q_{X}^{A}(1)}{Q_{X}^{B}(1)}\bigg)\bigg)k_{B}T$ i.e.
$\boxed{W=n\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T}$ (1)
This formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
The quantity (1) is maximised when Alice knows the prior $P_{X}$ and sets $Q_{X}^{A}=P_{X}$, in which case $W=D(P_{X}||Q_{X}^{B})k_{B}T$. This maximal extractable work from the system can be identified with the non-equilibrium free energy from thermodynamics up to a change in the base of the logarithm. In the particular case where Bob puts the partition in the middle, one recovers Szilard’s result $(1-H(P_{X}))\kappa_{B}T\ln{2}$ [22] in terms of entropy  [^1] It is useful to remember that for an isothermal process, the change in free energy for compression or expansion is given by $\Delta F=N\kappa_{B}T\ln(V_{f}/V_{i})$ where we have used the convention that work done on the gas is negative, and work done by the gas is positive..
The expression for the average extracted work by Alice is formally related to a formula appearing in the classical Kelly betting problem from information theory [18]: both involve a difference of relative entropies, and are equivalent up to an exponentiation. We develop this connection in detail in a companion paper  [^2] However, in Kelly’s gambling problem, wealth evolves multiplicatively, making the natural utility functions those of the CRRA type, as considered in [26].
It is also worth noting that Eq. (1) has also appeared in the work of [25] in the context of work extraction with an incorrect prior distribution. When Alice does not know the correct prior, our adversarial Szilard engine reduces to the formulation of [25] at the level of the ensemble average. However, here we are interested in the regime of finite-size $n$ without ensemble averages, and in the case where Alice knows the prior distribution. This will allow us to connect Eq. (1) to expected utility theory, economics, and Kelly betting, all of which assume a known distribution in the ideal case. We will thus assume that the prior distribution is known to Alice. Readers unfamiliar with Kelly betting may find it helpful to read the review in appendix A before proceeding to the next section.
III Utility maximisation: connection to expected utility formulations
In the previous section, we argued that the Kelly gambling problem and the adversarial Szilard engine work extraction problem were mathematically equivalent up to an exponentiation. The aim of this section is to connect stochastic thermodynamics to expected utility theory and resource theories using this observation.
Recall that varying levels of risk aversion are defined in expected utility theory by the willingness to compromise between amounts of wealth and uncertainty. In this sense, the optimal strategy for different agents depends on their particular level of risk aversion. A notable example connecting the Kelly utility function to other forms of risk aversion is given in [26], where the authors analyze Kelly gambling from the perspective of Constant Relative Risk Aversion (CRRA). CRRA describes a class of utility functions in which an individual’s relative risk aversion remains constant regardless of their level of wealth. For example, if a person has more wealth, they might risk a larger absolute amount while keeping the proportion of wealth they are willing to risk constant. This property makes CRRA particularly suitable for analyzing the Kelly paradigm, as it aligns with the proportional nature of Kelly betting, since it matches financial scenarios where wealth grows multiplicatively through compounding returns.
On the other hand, in thermodynamics and Szilard engines, the system responds to changes in fixed energy amounts. Losing a fixed amount of energy matters the same regardless of total system energy- this is captured in the economics literature by the CARA utility function. This utility function has already proven useful in other thermodynamic contexts [16]. The distinction between the two types of risk aversion reflects how thermodynamic systems respond to additive shocks, whilst financial systems experience multiplicative shocks (where gains and losses compound like interest).
We now explain how varying risk aversion plays a role in thermodynamic protocols. For readers without an economics background, let us begin by recalling that in expected utility theory, the most fundamental postulate is (called the expected utility hypothesis) states that rational individuals make decisions in order to maximise their expected utility, rather than expected value [27]. In this context, risk aversion is reflected in the utility function through the concavity - which reflects the fact that a risk-averse person might prefer a guaranteed $\mathdollar 100$ over a 50 per cent chance of winning $\mathdollar 250$, even though the latter has a higher expected value. Risk-seeking tendencies are defined analogously. Hence, strategies should be chosen to reflect these risk preferences.
The utility function we are considering in the thermodynamic context (CARA) is given by
$u_{r}(w_{x})=\frac{1}{r}(1-\exp(-rw_{x}))$ (2)
Here, the parameter $r$ reflects the risk preferences of the gambler
  • $r>0$: Risk-aversion (concave utility) - the agent prefers guaranteed work extraction over uncertain fluctuations
  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations
  • $r<0$: Risk-seeking behaviour (convex utility) - agents prefer higher amount of work extracted even if it means lower probability of success
As it is known from elementary thermodynamics [28], for a single round of work extraction in the generalised Szilard engine described above, the extracted work in terms of the outcome $x$ of the random variable $X$ can be expressed as
$w_{x}=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (3)
It follows using eq. 2 that the utility of outcome $x$ is
$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (4)
By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (5)
In appendix D, we show that for a given $r$, choice of $Q_{X}^{A}$ corresponds to
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (6)
Where we have introduced the notation $Q_{X}^{A,r}$ to indicate that the optimiser depends on the risk aversion of the gambler. Note that as $r$ goes to zero (risk neutral) $Q_{X}^{A,r}$ goes to $P_{X}$, and as $r$ goes to $\infty$ (extreme risk aversion) $Q_{X}^{A,r}$ goes to $Q_{X}^{B}$ (not moving partition, and hence no fluctuations at all), as expected.
In the following section, we connect the strategy $Q_{X}^{A,r}$ derived above the average work extracted by a gambler with a particular level of risk aversion, as well as to Renyi divergences (the generalised free energies of the resource-theoretic approach).
III.1 Risk aversion revisited
The aim of this section is to revisit risk aversion in the thermodynamic setting and to show explicitly how the economic notion of risk aversion helps provide the missing bridge between stochastic and resource-theoretic formulations of the second law.
Recall that, in the previous section, the utility maximisation problem of the adversarial Szilard engine led to optimal strategies for Alice of the form
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (7)
In our engine, this corresponds directly to Alice’s choice of partition placement; in economic terms, it reflects the unique optimal allocation for a CARA agent with parameter $r$. The parameter $r$ controls the risk aversion of the gambler [29] as follows:
  • $r>0$: Risk-aversion - the agent prefers guaranteed work extraction over uncertain fluctuations. Here risk aversion increases with increasing $r$.
  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations.
  • $r<0$: Risk-seeking behaviour - agents prefer higher amount of work extracted even if it means lower probability of success.
It is useful to evaluate the expected work extracted for a particular strategy. In appendix F, we show that for risk aversion level $r$, the expected work extracted $\mathbb{E}_{r}(Q_{X}^{A})$ is given by
$\boxed{\mathbb{E}_{r}[W]=\big(\alpha D(P_{X}||Q_{X}^{B})+(1-\alpha)D_{\alpha}(P_{X}||Q_{X}^{B})\big)k_{B}T}$ (8)
where we have included the subscript $\alpha=\frac{1}{1+r}$ to connect the level of risk aversion to the Renyi parameter and to indicate that the expected value of the extracted work depends on the particular strategy and level of risk aversion.
[…]
$\boxed{W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T}$ (9)
An analogous expression does not hold in the context of gambling, due to the fact that the utility function in that case is different. This result is complementary to the result of  [16] who quantified the dissipated fluctuating work $W_{CE}^{\mathrm{diss},\,r}$ in a thermal system which starts in equilibrium, and is driven out of it. This is the Crooks’ fluctuation relation setting. There, they find
$\beta W_{\mathrm{CE}}^{\mathrm{diss},\,r}=D_{1+r}\!\big(P_{\mathrm{F}}(w)\,\|\,P_{\mathrm{R}}(-w)\big),$ (10)
where $P_{F}(w)$ is the probability of the forward process and $P_{F}(-w)$ is the probability of the reverse process. Owing to the difference in initial starting states, it is unclear if these two expressions can be related, but it does allow and agent to decide between two different work extraction games in terms of expected utility.
It is useful to study the certainty equivalent and the expected work side by side. The expectation captures the stochastic-thermodynamic viewpoint of average entropy production, while the certainty equivalent aligns with the resource-theoretic notion of generalised free energies. Analysing their joint behaviour therefore clarifies how different levels of risk aversion interpolate between the two frameworks, and provides a rationality check on the agent’s strategy.
From the explanations above, we see that in the Szilard engine picture, the strategy for which the agent places the partition proportional to probability corresponds to risk neutrality $r=0$ in the economic formalism [29]. Indeed, in expected utility theory, a risk neutral individual is one for which the certainty equivalent equals the expected value of their gamble. For this risk attitude, both correspond to the thermodynamic non-equilibrium free energy.
Let us now focus on strategies corresponding to $r\geq 0$, i.e. risk aversion. Here the certainty equivalent $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})$ eq. 41 is non-negative, which follows directly from properties of the Renyi divergence in this regime [1]. Note also that since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the certainty equivalent decreases with increasing risk aversion, as expected. Since the Rényi divergence $D_{\alpha}$ is a non-decreasing function of its order $\alpha$ [30], and since $\alpha<1$ for $r>0$, it follows that $D_{\alpha}(P_{X}\|Q_{X}^{B})\leq D_{1}(P_{X}\|Q_{X}^{B})=D(P_{X}\|Q_{X}^{B})$. The weight $(1-\alpha)$ is positive, ensuring the second term is non-negative. However, because $D_{\alpha}(P_{X}\|Q_{X}^{B})$ is strictly less than $D(P_{X}\|Q_{X}^{B})$ for non-trivial distributions ($P_{X}\neq Q_{X}^{B}$), the weighted sum $\mathbb{E}[W]$ is also strictly less than the maximum $D(P_{X}\|Q_{X}^{B})$. This shortfall compared to the non-equilibrium free energy is the price the agent pays for certainty.
A subtlety arises for partition placements corresponding to risk-seeking behaviour ($r<0$), where the sign of both the certainty equivalent and the expected value may shift. Thermodynamically, risk-seeking strategies with negative expected value mark an operating point at which the agent aims to violate the second law: while rare fluctuations might yield temporary gains, they fail on average.
The rationality of the agent in such regimes is described by the joint behaviour of the certainty equivalent and the expected value in the following sense: if the certainty equivalent is more negative than the expected value of the gamble, the preferences remain consistent with expected utility theory. In this case, the agent is simply expressing a preference for the risky gamble which has a non-zero probability of a gain over a certain loss. This is consistent with risk-seeking behaviour.
However, risk-seeking behaviour becomes pathological if the certainty equivalent is more negative than the gamble’s worst possible outcome. In this scenario, the guaranteed loss prescribed by the certainty equivalent is strictly worse than every possible result of the gamble itself. This constitutes a violation of first-order stochastic dominance—a fundamental axiom of rational choice—as the dominated sure loss should never be chosen over the gamble that uniformly outperforms it [27, 31].
Hence, for risk-seeking behaviour, we must analyse the cases 1<r<0-1<r<0 and r<1r<-1 separately. For 1<r<0-1<r<0, the certainty equivalent remains positive and greater than the non-equilibrium free energy, reflecting the fact that the agent would only give up the gamble for a guaranteed amount greater than the non-equilibrium free energy.
For r<1r<-1, where the certainty equivalent flips sign, we must verify the stochastic dominance condition. In appendix G, we show that in our generalised Szilard engine, the condition for a violation of stochastic dominance in the regime where r<1r<-1 (the regime for which the sign of the certainty equivalent flips) is given by
$D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$ (11)
However, the right hand side is equal to $D_{-\infty}(P_{X}||Q_{X}^{B})$, so the condition reduces to $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})<D_{-\infty}(P_{X}||Q_{X}^{B})$. But since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the first-order stochastic dominance condition is never violated.
IV Thermodynamic risk aversion
In the section II, we showed that the non-equilibrium free energy from thermodynamics can be understood as the maximum amount of work that can be extracted in the presence of an adversary at the level of the ensemble average. The aim of this section is to analyse the same ensemble from a finite-size regime perspective using tools developed in [17], and to prove that work extraction in the finite-size regime corresponds to the maximisation of a utility function as described in section III.
Recall first that for a single round of work extraction, the extracted work in terms of the outcome $x$ of the random variable can be expressed as
$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T.$ (12)
Again, this formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
It follows that for $n$ rounds, the work extracted is given by
$W_{n}=\sum_{i=1}^{n}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (13)
by grouping terms with the same outcome, and denoting by $N_{x}$ the number of times each outcome occurs, this can be expressed as
$W_{n}=\sum_{x}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}k_{B}T$ (14)
It is useful to define the type $\lambda_{x^{n}}$ of the sequence $x^{n}$. Often called its empirical distribution, the type is defined as $\lambda_{x^{n}}(x)=N_{x}/n$. Rewriting the equation above in terms of the type, we obtain an equation analogous to eq. 1 in the finite-size regime:
$\boxed{W_{n}=n(D(\lambda_{x^{n}}||Q^{B}_{X})-D(\lambda_{x^{n}}||Q^{A}_{X}))k_{B}T}$ (15)
Hence, by placing the partition at position $Q_{X}^{A}=\lambda_{x^{n}}$ (when Alice’s allocation matches the type of the observed sequence), the work extracted is given by $n(D(\lambda_{x^{n}}||Q^{B}_{X}))k_{B}T$. In the finite-size regime, the work extraction problem reduces to what is essentially a decision-theoretic problem of “guessing the type” of the empirical sequence $\lambda_{x^{n}}$ that will be realised. In concurrent work [17], we show that each possible type corresponds to a particular fluctuation, with probability decaying exponentially up to sub-exponential factors as $P_{X}(\lambda_{x^{n}})\doteq\exp(-nD(\lambda_{x^{n}}\|P_{X}))$, while the potential work reward scales as $nD(\lambda_{x^{n}}\|Q_{X}^{B})$.
This creates a fundamental risk-reward trade-off: Alice must select her strategy $Q_{X}^{A}$ to balance the probability of success against the amount of work extracted when successful. In appendix C, we show that the optimal strategy that maximises work extraction for a given risk tolerance (minimum success probability $\epsilon$) takes the form of an exponential family interpolating between $P_{X}$ and $Q_{X}^{B}$:
$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (16)
where the parameter $\mu$ is determined by the risk constraint. The corresponding work bound is given in terms of a Rényi divergence:
$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (17)
The detailed derivation using the method of types [32] is based on a treatment of Kelly gambling in the finite-size regime developed in concurrent work [17]. This approach reveals that all Rényi divergences $D_{\alpha}$ acquire an operational interpretation in terms of work extraction for different risk tolerances. Previous work [20] provided operational interpretation for min and max entropies as bounds in finite-size work extraction. Here, we find similar operational interpretation for all Renyi divergences. Crucially, we also provide the work extraction strategy given the agent’s constraints. In addition, this result is identical to the expected utility formulation of the problem after identifying
$\frac{r}{1+r}=1-\frac{1}{1+r}$ (18)
with $\mu$ in eq. 16.
V Conclusions
We have shown that the problem of finite-size work extraction can be understood within a resource-theoretic formulation of adversarial gambling, in which an agent’s strategic choice plays the role of a utility-maximizing gamble against a thermodynamic adversary. This perspective reveals that the relevant notion of risk aversion in thermodynamics is governed by CARA utilities, reflecting the additivity of work increments, in contrast to the CRRA utilities that arise in the multiplicative wealth growth of Kelly betting. This structural difference explains why, in the thermodynamic setting, the certainty equivalent of extracted work coincides with Rényi divergences, thereby grounding the generalized second laws in an operational principle of expected-utility maximization.
The identification of the relevant utility function shows that incorporating risk aversion into work-extraction protocols makes both stochastic sensitivity to fluctuations and the generalized free energies of thermodynamics emerge from a single principle of decision theory. In fact, the certainty equivalents defined by CARA utilities coincide exactly with Rényi divergences, demonstrating that the two modern perspectives—the stochastic description of fluctuating work and the resource-theoretic hierarchy of free energies—are not separate constraints but two ways of expressing the same decision-theoretic principles. While the generalised free energies found in the context of second laws are given in terms of the Renyi-divergence from the initial state $P_{X}$ to the equilibrium state at inverse temperature $\beta$[8], here we found that the relevant Renyi-divergences is from the state $P_{X}$ to the out of equilibrium state $Q_{X}^{B}$. Understanding the relationships between these, is an interesting open question.
Acknowledgements
MA thanks Ian Ford and Paul Skrzypczyk for useful discussions. MA was supported by the Mexican National Council of Science and Technology (CONACYT), EPSRC, the Simons Foundation It from Qubit Network, and SNF Quantum Flagship Replacement Scheme (Grant No. 215933). PhF is supported by the project FOR 2724 of the Deutsche Forschungsgemeinschaft (DFG). TS is supported by by JST ERATO Grant Number JPMJER2302, Japan. JO was supported by the Simons Foundation, It from Qubit Network, and EPSRC.
Appendix A Review of Kelly betting
The Kelly [18] betting scheme is an adversarial set-up in which the gambler, Alice, allocates fractions of her wealth to the possible outcomes of a horse race, always distributing a non-zero fraction to each outcome (in order to avoid ever being completely broke). The adversary Bob, also sets odds for each possible outcome.
Suppose that the horse race is described by the random variable $X$. Bob, sets the odds $o_{x}$ to outcome $x$, and Alice allocates the fraction $f_{x}$ of her wealth to outcome $x$. With this convention for describing odds, after a round of the game, Alice’s initial wealth $W_{i}$ is multiplied by the factor $f_{x}o_{x}$:
$W_{1}=f_{x}o_{x}W_{i}$ (19)
Assuming that Alice re-invests whatever wealth she has after the previous round, her wealth evolves recursively, and at round $N$ it obeys the equation
$W_{n}=f_{x}o_{x}W_{n-1}$ (20)
Since Alice reinvests her wealth after each round, her wealth after $n$ rounds can be expressed as a product of the wealth multipliers for each outcome:
$\frac{W_{n}}{W_{i}}=\prod_{x}(f_{x}o_{x})^{N_{x}}$ (21)
where $N_{x}$ is the number of times that the outcome $x$ occurred.
Since the fraction by which the wealth is multiplied after every round is a random variable, it is customary to write $Q_{X}^{B}(x)=o_{x}^{-1}$ and $Q^{A}_{X}(x)=f_{x}$. Here, the subscript is there to $X$ indicate that the ratio $\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}$ is a random variable, whilst the superscripts $A$ and $B$ indicate that the fraction is chosen by Alice (resp. Bob). Altogether, the notation helps to stress the fact that even though the bets and odds are the same at every round of gambling, the amount of money made by Alice is a random variable. The expression above now becomes
$W_{n}=\prod_{x}\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}W_{i}$ (22)
In the limit as $n>>1$, the ratio of Alice’s initial wealth $W_{i}$ to her final wealth $W_{F}$ satisfies:
$\frac{W_{F}}{W_{i}}=\exp(n(D(P_{X}||Q_{X}^{B})-D(P_{X}||Q_{X}^{A})))$ (23)
Here, $D(P_{X}\|Q_{X})$ is the relative entropy (or Kullback-Leibler divergence). Though not a true distance metric, it is often interpreted in information theory as measuring the discrepancy between probability distributions. Since relative entropy is non-negative, Alice’s optimal strategy is to set $Q_{X}^{A}=P_{X}$, ensuring that her wealth grows at the maximum possible rate. The expression above is the well-known result by Kelly [18].
This multiplicative growth of wealth under Kelly gambling contrasts with the additive accumulation of work in the thermodynamic engine, fundamentally shaping the respective utility functions (CRRA vs. CARA) that describe rational behavior in each domain.
Appendix B General protocol for adversarial work extraction
In this section, we derive the work formula for a general adversarial engine. Let $x$ be a microscopic state that is not necessarily binary. Suppose that the initial distribution is given by $P_{X}(x)$ and the initial energy level is given by $E_{X}^{B}(x)$. We relate the energy level to Bob’s distribution by $Q_{X}^{B}(x)=e^{-E_{X}^{B}(x)}/Z_{X}^{B}$ with $Z_{X}^{B}=\sum_{x}e^{-E_{X}^{B}(x)}$ being the normalization factor (partition function). Note that we set $k_{B}T$ to unity throughout this section.
Let us first remember the optimal work extraction protocol at the level of ensemble average: (i) Alice quenches the energy level to $E_{X}(x)$ that is defined through $P_{X}(x)=e^{-E_{X}(x)}/Z_{X}$ with $Z_{X}=\sum_{x}e^{-E_{X}(x)}$; (ii) Alice moves the energy level from $E_{X}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally. It is well known that Alice extracts the work $D(P_{X}\|Q_{X}^{B})k_{B}T$ on average by this protocol [1, 3].
Now we turn to the adversarial scenario. The protocol that Alice performs is: (i) Alice quenches the energy level to $E_{X}^{A}(x)$ that is defined through $Q_{X}^{A}(x)=e^{-E_{X}^{A}(x)}/Z_{X}^{A}$ with $Z_{X}^{A}=\sum_{x}e^{-E_{X}^{A}(x)}$ for her choice of $Q_{X}^{A}$; (i’) Alice lets the system thermalized; (ii) Alice moves the energy level from $E_{X}^{A}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally. The extracted work in step (i) is given by $E_{X}^{B}(x)-E_{X}^{A}(x)$, while in step (ii) $\ln Z_{X}^{B}-\ln Z_{X}^{A}$ (the change in the equilibrium free energy). We thus obtain
$w_{X}(x)=E_{X}^{B}(x)-E_{X}^{A}(x)+\ln Z_{X}^{B}-\ln Z_{X}^{A}=\ln\left(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\right),$ (24)
which reproduces Eq. (12). By taking the average with respect to $P_{X}$, Eq. (1) is also reproduced.
Appendix C Work extraction and the method of types
This appendix summarizes the application of the method of types to the adversarial Szilard engine, following the general approach detailed in concurrent work [17]. The method of types [32] provides a powerful framework for analyzing finite-size fluctuations in information-theoretic tasks.
For a sequence of $n$ rounds, the empirical distribution (type) $\lambda_{x^{n}}$ occurs with probability approximately $\exp(-nD(\lambda_{x^{n}}\|P_{X}))$. The work extracted when the type is $\lambda_{x^{n}}$ is $nD(\lambda_{x^{n}}\|Q_{X}^{B})k_{B}T$ when Alice sets $Q_{X}^{A}=\lambda_{x^{n}}$.
The optimization problem becomes: maximize the reward $D(\lambda_{x^{n}}\|Q_{X}^{B})$ subject to the constraint that the probability of success exceeds $\epsilon$, i.e., $D(\lambda_{x^{n}}\|P_{X})\leq\frac{1}{n}\ln(1/\epsilon)$. Using Lagrange multipliers, one obtains the family of optimal strategies:
$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (25)
where $\mu\in[0,1]$ is determined by the constraint. Substituting back gives the work bound:
$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (26)
For a comprehensive treatment of the method of types in adversarial scenarios, including the connection to hypothesis testing and utility theory, we refer to [17, 21].
Appendix D Maximisation of the utility function
In this section, we compute the maximiser of the utility function for risk parameter $r$. Recall now that the extracted work from the adversarial Szilard engine for a single round can be described by the random variable
$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (27)
It follows using eq. 2 that the utility of outcome $x$ is
$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (28)
By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (29)
Note that the first term is a constant, so we need only focus on the second term. For $r>0$ (risk-aversion), the second term is negative, whilst if $r<0$ (risk-seeking), the second term is positive. This means that for a risk-averse individual, we minimise
$\min_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (30)
which is equivalent to minimising
$\min_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (31)
whilst for a risk-seeking individual, we maximise
$\max_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (32)
which is equivalent to maximising
$\max_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (33)
In order to find the minimiser (resp. maximiser), we will use a lemma proved in [26], which states that
$\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})=-r\big(D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})\big)$ (34)
where
$G_{X}^{r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (35)
The maximiser is therefore
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (36)
Appendix E Calculation of the certainty equivalent
At the beginning of this section, we said that two equivalent ways of characterising risk preferences were through the concavity/convexity of the utility function and the certainty equivalent. As shown in [16], the certainty equivalent has a clear information-theoretic interpretation in a thermodynamic context. We will also find it useful here to calculate the certainty equivalent of work for each value of the risk aversion parameter $r$ below.
Recall that, for an agent exhibiting risk aversion, the certainty equivalent is the guaranteed amount of money that the individual would accept instead of taking a risky bet with different payoff. We then say that an individual with utility function $u(.)$ is risk averse if and only if their certainty equivalent is lower than the expected value of the risky lottery. The mathematical definition of the certainty equivalent is the number $W_{CE}\in\mathbb{R}$ such that $u_{r}(W_{CE})=E(u_{r}(w_{X}(x)))$, where $E$ denotes expected value. Since the utility of outcome $x$ is
$u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (37)
we must calculate the number $W_{CE}$ $u_{r}(W_{CE})=E(u_{r}(x))$, where $E$ denotes expected value, we therefore calculate
$W_{CE}=u_{r}^{-1}\big(E(u_{r}(w))\big)\\ =u_{r}^{-1}\bigg(E\bigg(\frac{1}{r}(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =u_{r}^{-1}\bigg(\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (38)
where in the last line we have used that the inverse function of $u_{r}(x)$ is $u_{r}^{-1}(x)=-\frac{1}{r}\ln(1-rx)$. Since
$W_{CE}=-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (39)
which equals
$D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})$ (40)
by the result of [26], we finally conclude that
$W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T$ (41)
when the expression for $Q_{X}^{A,r}$ is substituted into the expression for the certainty equivalent.
Since $D_{\alpha}<D_{\beta}$ for $\alpha<\beta$, one sees that for risk-averse (resp. risk-seeking) individuals, the certainty equivalent decreases (resp. increases) with increasing (resp. decreasing) risk aversion, as expected.
In addition, the work certainty equivalent is smaller than $D_{1}$ (the non-equilibrium free energy) for a risk averse player, and equal to $D_{1}$ for a risk neutral player. In particular, for an extremely risk averse player ($r\to\infty$), the certainty equivalent tends to $D_{0}(P_{X}||Q_{X}^{B})$, which becomes zero when both distributions have the same support.
Appendix F Extracted work as a function of strategy
In section II, we showed that for a given strategy of Alice $Q_{X}^{A}$, the average work extracted in the adversarial Szilard engine is given by
$\mathbb{E}[W]=\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T$ (42)
Our goal is to express this for the optimal strategy $Q_{X}^{A}$ which from now on we write as $Q^{A,r}_{X}$, to emphasise dependence on the risk parameter of the CARA utility function. We substitute the explicit form of the optimal strategy into the second term.
The optimal strategy for a given risk parameter $r$ is:
$Q^{A,r}X(x)=\frac{P_{X}(x)^{\frac{1}{1+r}}Q^{B}_{X}(x)^{\frac{r}{1+r}}}{Z},$ (43)
where
$Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q^{B}_{X}(x^{\prime})^{\frac{r}{1+r}}$ (44)
Let $\alpha=\frac{1}{1+r}$ for notational clarity. Since the first divergence is independent of the strategy $Q^{A,r}_{X}$, we compute the second divergence term, $D(P_{X}||Q^{A,r}_{X})$:
$D(P_{X}||Q^{A,r}_{X})$ $=\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{A,r}_{X}(x)}$ (45)
$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\ln\left(\frac{P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}}{Z}\right)\right]$ (46)
$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\alpha\ln P_{X}(x)-(1-\alpha)\ln Q^{B}_{X}(x)+\ln Z\right]$ (47)
$=(1-\alpha)\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{B}_{X}(x)}+\ln Z$ (48)
$=(1-\alpha)D(P_{X}||Q^{B}_{X})+\ln Z(\alpha)$ (49)
The normalization constant $Z$ is related to the sum that defines the Rényi divergence. Recall its definition:
$D_{\alpha}(P_{X}||Q^{B}_{X})=\frac{1}{\alpha-1}\ln\sum_{x}P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}=\frac{1}{\alpha-1}\ln Z$ (50)
So that $\ln Z$:
$\ln Z(\alpha)=(\alpha-1)D_{\alpha}(P_{X}||Q^{B}_{X})$ (51)
Substituting (51) back into (45):
$D(P_{X}||Q^{A,r}X)$ $=(1-\alpha)D(P_{X}||Q^{B}_{X})+(\alpha-1)D\alpha(P_{X}||Q^{B}_{X})$ (52)
$=(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]$ (53)
Finally, we substitute this result back into the original expression for expected work (42):
$\mathbb{E})[W]=D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A,r}X)\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})\\ =\alpha D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})$ (54)
Thus, the expected work extracted for the optimal strategy corresponding to risk parameter $r$ is:
$\boxed{\mathbb{E}[W]=\alpha D(P_{X}||Q^{B}_{X})k_{B}T+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})k_{B}T,\quad\text{where}\quad\alpha=\frac{1}{1+r}}$ (55)
Appendix G Conditions for rationality for r<1r<-1
In section eq. 41 we showed that the certainty equivalent for work extraction for a rational agent with CARA utility function is a Rényi divergence:
$W_{CE}=D_{\alpha}(P_{X}||Q_{X}^{B})k_{B}T\quad\text{where}\quad\alpha=\frac{1}{1+r}$ (56)
The work extracted in a specific outcome $x$ is:
$W(x)=\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)k_{B}T$
where $Q_{X}^{A,r}$ is the optimal strategy for a given value of the risk parameter $r$.
The violation of rationality (First-Order Stochastic Dominance) occurs if the certainty equivalent is more negative than the worst-case work output:
$W_{CE}<\min_{x}W(x)$
Substituting in the expression for the certainty equivalent, we get:
$D_{\alpha}(P_{X}||Q_{X}^{B})<\min_{x}\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)$
This is the general condition. We can now use the specific form of $Q_{X}^{A,r}$ to simplify the right-hand side. Recall:
$Q_{X}^{A,r}(x)=\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}}$
Defining $Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}$ as the normalisation constant and substituting into the work expression:
$W(x)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{Z(\alpha)Q_{X}^{B}(x)}\right)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{-\alpha}}{Z(\alpha)}\right)=\alpha\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
To find the minimum value of this expression, note that the term $\ln Z(\alpha)$ is independent of $x$. Therefore:
$\min_{x}W(x)=\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
The normalisation constant $\ln Z(\alpha)$ is directly related to the Rényi divergence by:
$D_{\alpha}(P_{X}||Q_{X}^{B})=\frac{1}{\alpha-1}\ln Z(\alpha)$
Now, we can substitute back into our violation condition (Eq. 1):
$D_{\alpha}(P_{X}||Q_{X}^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
Substitute $D_{\alpha}$ and rearrange:
$\frac{1}{\alpha-1}\ln Z(\alpha)<\alpha\cdot M-\ln Z(\alpha)\quad\text{where}\quad M=\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$
$\ln Z(\alpha)\left(\frac{1}{\alpha-1}+1\right)<\alpha M$
$\ln Z(\alpha)\left(\frac{1+(\alpha-1)}{\alpha-1}\right)<\alpha M$
$\frac{\ln Z(\alpha)\cdot\alpha}{\alpha-1}<\alpha M$
For $\alpha<0$ (which is true for r<1r<-1, the case we are interested in), dividing both sides by $\alpha$ flips the sign of the inequality:
$\frac{\ln Z(\alpha)}{\alpha-1}>M$
Since the left-hand side is exactly the Rényi divergence $D_{\alpha}(P_{X}||Q_{X}^{B})$, the condition for a violation of rationality simplifies to:
$\boxed{D_{\alpha}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)}$
since this never holds, rationality is never violated.

### v2

@@
 Traditional frameworks struggle to capture how strategic decision-making under uncertainty — particularly an agent’s tolerance for risk — determines the trade-off between extractable work and probability of success in finite-scale experiments.
-Here, we develop a framework for non-equilibrium thermodynamics based on adversarial resource theories, in which work extraction is modelled as an adversarial game for an agent extracting work.
-Within this perspective, we recast the Szilard engine as a game isomorphic to Kelly gambling, an information-theoretic model of optimal betting under uncertainty — but with a thermodynamic utility function.
-Extending the framework to finite-size regimes, we apply a risk-reward trade-off to find an interpretation of the Renyi-divergences, in terms of extractable work for a given failure probability.
+Here, we develop a framework for nonequilibrium thermodynamics based on adversarial resource theories, in which work extraction is modeled as an adversarial game for an agent extracting work.
+Within this perspective, we consider a Szilard-type engine as a game isomorphic to Kelly gambling, an information-theoretic model of optimal betting under uncertainty — but with a thermodynamic utility function.
+Extending the framework to finite-size regimes, we apply a risk–reward trade-off to find an interpretation of the Rényi-divergences, in terms of extractable work for a given failure probability.
 By incorporating risk sensitivity via utility functions, we show that the guaranteed amount of work a rational agent would accept instead of undertaking a risky protocol is given by a Rényi divergence.
@@
 Traditionally, for large systems at equilibrium, the second law can be understood in two complementary ways: either as a constraint on average entropy production (or, equivalently, as a decrease in free energy) or as a rule governing which state transitions are physically possible.
-In recent decades, advances in nanotechnology and biophysics have made the study of non-equilibrium and small-scale thermodynamics increasingly important.
-In these regimes—where the thermodynamic limit no longer applies and systems contain only finitely many interacting particles—ensemble averages fail and fluctuations dominate.
-Deterministic quantities such as free energy must be replaced by probabilistic, protocol-dependent notions, leading to tradeoffs between extractable work and probability of success.
-The breakdown of determinism has motivated two main approaches to understand the second law of thermodynamics in regimes where standard assumptions break down: stochastic thermodynamics, and the resource-theoretic approach (for a review see [1]).
-Stochastic thermodynamics[2, 3] embraces the inherent randomness of small, out-of-equilibrium systems, by treating work as a fluctuating quantity.
+In recent decades, advances in nanotechnology and biophysics have made the study of nonequilibrium and small-scale thermodynamics increasingly important.
+In these regimes where the thermodynamic limit no longer applies and systems contain only finitely many interacting particles, ensemble averages fail and fluctuations dominate.
+Deterministic quantities such as free energy must be replaced by probabilistic, protocol-dependent notions, leading to trade-offs between extractable work and probability of success.
+The breakdown of determinism has motivated two main approaches to understand the second law of thermodynamics in regimes where standard assumptions break down: stochastic thermodynamics, and the resource-theoretic approach (for a review, see Ref.
+Sagawa (2022)).
+Stochastic thermodynamics Seifert (2012); Parrondo et al.
+(2015) embraces the inherent randomness of small, out-of-equilibrium systems by treating work as a fluctuating quantity.
 It describes how entropy on average must increase, while accounting for rare transient decreases.
-On the other hand, the resource-theoretic approach[4, 5, 6] reformulates the second law as a constraint on possible state transformations.
-It shifts the focus from averages to operational possibilities, yielding generalized second laws (expressed through Renyi divergences) that quantify how far a state is from thermal equilibrium[7, 8].
-While both frameworks are profoundly successful, and their views on the second law can be related[9], neither provides a complete, operational prescription for how an agent’s strategic choices—specifically, their tolerance for risk—directly determine the trade-off between the amount of work they can extract and the probability of successfully obtaining it in a single, finite-scale experiment.
-In this work, we bridge this gap by treating work extraction from the lens of expected utility theory and decision theory [10].
-There have been seminal works forging connections between stochastic thermodynamics and gambling  [11, 12, 13, 14, 15, 16].
-One strand of research has implemented specific gambling strategies directly within thermodynamic protocols, as seen in the work extraction experiments of [13, 14].
-Concurrently, [12] established formal links between gambling, work extraction, and information flows from an information-thermodynamic perspective.
-In a separate approach, [16] applied expected utility theory to evaluate thermodynamic processes as lotteries with pre-defined work payoffs.
+On the other hand, the resource-theoretic approach Janzing et al.
+(2000); Horodecki et al.
+(2003); Horodecki and Oppenheim (2013) reformulates the second law as a constraint on possible state transformations.
+It shifts the focus from averages to operational possibilities, yielding generalized second laws — expressed through Rényi divergences — that quantify how far a state is from thermal equilibrium Ruch and Mead (1976); Brandão et al.
+(2015).
+While both frameworks are profoundly successful, and their views on the second law can be related Alhambra et al.
+(2016), neither provides a complete, operational prescription for how an agent’s strategic choices — specifically, their tolerance for risk — directly determine the trade-off between the amount of work they can extract and the probability of successfully obtaining it in a single, finite-scale experiment.
+In this work, we bridge this gap by treating work extraction from the lens of expected utility theory and decision theory Von Neumann and Morgenstern (1944).
+There have been seminal works forging connections between stochastic thermodynamics and gambling Hirono and Hidaka (2015); Ito (2016); Manzano et al.
+(2020); Tohme et al.
+(2023); Tirone et al.
+(2021); Ducuara et al.
+(2023).
+One strand of research has implemented specific gambling strategies directly within thermodynamic protocols, as seen in the work extraction experiments of Manzano et al.
+(2020); Tohme et al.
+(2023).
+Concurrently, Ref.
+Ito (2016) established formal links between gambling, work extraction, and information flows from an information-thermodynamic perspective.
+In a separate approach, Ref.
+Ducuara et al.
+(2023) applied expected utility theory to evaluate thermodynamic processes as lotteries with predefined work payoffs.
 While these works successfully import concepts from one field to the other, our framework provides a deeper analogy between them and reveals a decision-theoretic aspect of thermodynamics.
 By formally modeling work extraction as an adversarial game implemented by a Szilard-type engine, we demonstrate that generalized free energies emerge from the necessary constraints on rational, strategic play.
-We develop a framework based on adversarial resource theories [17], in which an agent must choose a work-extraction protocol while contending with constraints initially imposed on the engine.
+We develop a framework based on adversarial resource theories Arcos et al.
+(2025), in which an agent must choose a work-extraction protocol while contending with constraints initially imposed on the engine.
 This perspective casts thermodynamics as a decision-theoretic problem, where the optimal strategy is determined by the agent’s sensitivity to fluctuations.
-The mathematical structure of this game is formally analogous to the Kelly betting problem from information theory [18], a connection we develop in a companion paper [17].
+The mathematical structure of this game is formally analogous to the Kelly betting problem from information theory Kelly (1956), a connection we develop in a companion paper Arcos et al.
+(2025).
 However, a crucial distinction arises from the physical context: while financial wealth in Kelly gambling grows multiplicatively, thermodynamic work is an additive quantity.
@@
 The paper is structured as follows.
-In section, II, we formulate a Szilard-type engine as an adversarial setup between an agent (Alice) extracting the work and another agent (Bob) preparing initial constraints, and derive the average extractable work and its connection to the non-equilibrium free energy.
-In section III, we analyse the risk-reward trade-off through the lens of decision theory, and in particular expected utility theory, where we identify Constant Absolute Risk Aversion (CARA) as the relevant utility class for thermodynamics.
-This framework allows us to compute both the certainty equivalent (the guaranteed amount of work a risk-sensitive individual would accept instead of undertaking a risky extraction protocol) as Eq (9), and the expected value of the work extracted by a rational individual, Eq (8) for any level of risk aversion.
-We find that both are parametrized by Renyi divergences.
-We explain why the former result has no analogue in gambling, while the latter result is related to the gambling result derived in [17].
+In Section II, we formulate a Szilard-type engine as an adversarial setup between an agent (Alice) extracting the work and another agent (Bob) preparing initial constraints, and derive the average extractable work and its connection to the nonequilibrium free energy.
+In Section III, we analyze the risk–reward trade-off through the lens of decision theory, and in particular expected utility theory, where we identify Constant Absolute Risk Aversion (CARA) as the relevant utility class for thermodynamics.
+This framework allows us to compute both the certainty equivalent (the guaranteed amount of work a risk-sensitive individual would accept instead of undertaking a risky extraction protocol) as Eq.
+(11), and the expected value of the work extracted by a rational individual, Eq.
+(9) for any level of risk aversion.
+We find that both are parametrized by Rényi divergences.
+We explain why the former result has no analogue in gambling, while the latter result is related to the gambling result derived in Arcos et al.
+(2025).
 In Section IV, we approach the problem from an information-theoretic viewpoint on the finite-size regime, where fluctuations dominate.
-Here, we apply results from concurrent work [19, 17] to show that the work extraction problem reduces to a decision-theoretic problem, introducing a fundamental risk-reward trade-off in which the work extracted if a strategy is successful can be bounded in terms of Renyi divergences $D_{\alpha}$.
-This gives an operational interpretation to $D_{\alpha}$ for each individual $\alpha$, in terms of the minimal work extraction given some risk tolerance.
-This extends the result of  [20] who gave an interpretation of the Renyi divergences $D_{0}$ and $D_{\infty}$ in terms of bounds on extracting work in the case of an extremely risk averse or risk seeking agent.
-In both the expected utility approach and the information-theoretic approach, we give identical and explicit strategies for achieving work extraction given some risk tolerance via Eq (7).
+Here, we apply results from concurrent work Arcos Enriquez (2025); Arcos et al.
+(2025) to show that the work extraction problem reduces to a decision-theoretic problem, introducing a fundamental risk–reward trade-off in which the work extracted if a strategy is successful can be bounded in terms of Rényi divergences $D_{\alpha}$.
+This gives an operational interpretation to $D_{\alpha}$ for each value of $\alpha$, in terms of the minimal work extraction given some risk tolerance.
+This extends the result of Dahlsten et al.
+(2011) who gave an interpretation of the Rényi divergences $D_{0}$ and $D_{\infty}$ in terms of bounds on extracting work in the case of an extremely risk-averse or risk-seeking agent.
+In both the expected utility approach and the information-theoretic approach, we give identical and explicit strategies for achieving work extraction given some risk tolerance, via Eq.
+(8).
 Together, all these results provide a unified operational interpretation that bridges stochastic and resource-theoretic views of the second law.
@@
 II The adversarial Szilard engine
-We begin by formalising a thermodynamic work extraction problem as a set-up between adversaries, mirroring the structure of Kelly betting, which we review in appendix A.
+We begin by formalizing a thermodynamic work extraction problem as a setup between adversaries, mirroring the structure of Kelly betting, which we review in Appendix A.
 Consider three players: Bob, who sets initial constraints; Alice, who optimizes work extraction (analogous to a gambler allocating bets), and Charlie, a referee who enforces randomness.
-We consider an empty box of volume $V$ which Bob divides into two parts by placing a partition at some position $Q_{B}$.
-The referee, Charlie, then samples from a binary probability distribution $P_{X}(x)$ and places a molecule on the left or right hand side of the box according to the outcome which Alice and Bob do not know.
-Alice proceeds to extract work by performing isothermal compressions and expansions on the box, moving the partition to a final position $Q_{A}$ of her choice.
+We consider an empty box of volume $V$ which Bob divides into two parts by placing a partition at some position $Q^{B}$.
+The referee, Charlie, then samples from a binary probability distribution $P(x)$ and places a molecule on the left or right-hand side of the box according to the outcome which Alice and Bob do not know.
+Here, $x=0$ (11) represents that the molecule is in the left (right) side.
+Alice proceeds to extract work by performing isothermal compressions and expansions on the box, moving the partition to a final position $Q^{A}$ of her choice.
 The process is illustrated in Fig.
 1.
-(a) Bob places the partition at some initial position $Q^{B}$
-(b) Alice chooses a position $Q^{A}$ to which she moves the partition
-(c) When the molecule is on the left hand side of the box, Alice extracts work by moving it to her desired position
-(d) When the molecule is on the right hand side of the box, Alice must do work to move the partition to the position of her choice
+Since this setup resembles the Szilard engine Szilard (1929), we name it the adversarial Szilard engine.
+However, we emphasize that our setup does not involve measurement and feedback by an external “demon”, in contrast to the original Szilard’s setup.
+(a) Bob places the partition at some initial position $Q^{B}$.
+(b) Alice chooses a position $Q^{A}$ to which she moves the partition.
+(c) When the molecule is on the left side of the box, Alice extracts work by moving it to her desired position.
+(d) When the molecule is on the right side of the box, Alice must do work to move the partition to the position of her choice.
 Figure 1: Schematic of the adversarial Szilard engine.
-Without loss of generality, assume that Bob places the partition closer to the leftmost edge of the box, so that the volume on the left hand side can be expressed as $Q_{B}\textit{l}A$, and that Alice moves her partition to a position so that the volume on the left hand side can be expressed as $Q_{A}\textit{l}A$ in terms of fractions $Q_{A}\in(0,1)$ and $Q_{B}\in(0,1)$ of the total length $l$ of the box.
-The work extracted by Alice depends on the molecule’s position:
-  • When $x=0$ (left-hand side): $w_{0}=k_{B}T\ln(Q^{A}/Q^{B})$
-  • When $x=1$ (right-hand side): $w_{1}=k_{B}T\ln\left(\frac{1-Q^{A}}{1-Q^{B}}\right)$
-To maintain consistency with the notation of the literature [21], we write $Q^{B}=Q_{X}^{B}(0)$ and $1-Q^{B}=Q_{X}^{B}(1)$, and similarly for Alice.
+Without loss of generality, assume that Bob places the partition closer to the leftmost edge of the box, so that the volume on the left-hand side can be expressed as $Q^{B}V$, and that Alice moves her partition to a position so that the volume on the left-hand side can be expressed as $Q^{A}V$ in terms of fractions $0<Q^{A}<1$ and $0<Q^{B}<1$ of the total length of the box.
+The work extracted by Alice depends on the molecule’s position $x=0,1$:
+$w(0)=k_{B}T\ln\left(\frac{Q^{A}}{Q^{B}}\right);\ \ w(1)=k_{B}T\ln\left(\frac{1-Q^{A}}{1-Q^{B}}\right).$ (1)
+To maintain consistency with the notation of the literature Cover and Thomas (2006), we write $Q^{B}(0):=Q^{B}$ and $Q^{B}(1):=1-Q^{B}$, and similarly for Alice.
 With this notation, the amount of work extracted (which is a random variable) can be written in terms of the positions to which Alice and Bob move the partition (which are not a random variable).
-When the process is repeated a large number $n$ of times, the average amount of work $W$ extracted by Alice is given by $W=n\bigg(P_{X}(0)\ln{\bigg(\frac{Q_{X}^{A}(0)}{Q_{X}^{B}(0)}}\bigg)+P_{X}(1)\ln\bigg(\frac{Q_{X}^{A}(1)}{Q_{X}^{B}(1)}\bigg)\bigg)k_{B}T$ i.e.
-$\boxed{W=n\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T}$ (1)
-This formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
-The quantity (1) is maximised when Alice knows the prior $P_{X}$ and sets $Q_{X}^{A}=P_{X}$, in which case $W=D(P_{X}||Q_{X}^{B})k_{B}T$.
-This maximal extractable work from the system can be identified with the non-equilibrium free energy from thermodynamics up to a change in the base of the logarithm.
-In the particular case where Bob puts the partition in the middle, one recovers Szilard’s result $(1-H(P_{X}))\kappa_{B}T\ln{2}$ [22] in terms of entropy  [^1] It is useful to remember that for an isothermal process, the change in free energy for compression or expansion is given by $\Delta F=N\kappa_{B}T\ln(V_{f}/V_{i})$ where we have used the convention that work done on the gas is negative, and work done by the gas is positive..
-The expression for the average extracted work by Alice is formally related to a formula appearing in the classical Kelly betting problem from information theory [18]: both involve a difference of relative entropies, and are equivalent up to an exponentiation.
-We develop this connection in detail in a companion paper  [^2] However, in Kelly’s gambling problem, wealth evolves multiplicatively, making the natural utility functions those of the CRRA type, as considered in [26].
+The average amount of work $W$ extracted by Alice is given by
+$W=k_{B}T\bigg(P(0)\ln\bigg(\frac{Q^{A}(0)}{Q^{B}(0)}\bigg)+P(1)\ln\bigg(\frac{Q^{A}(1)}{Q^{B}(1)}\bigg)\bigg),$ (2)
+or equivalently,
+$\boxed{W=k_{B}T\big(D(P||Q^{B})-D(P||Q^{A})\big)}$ (3)
+where $D(\cdot\|\cdot)$ is the standard relative entropy (the Kullback-Leibler divergence).
+This formula is true not only for the above binary adversarial Szilard engine (Fig.
+1), but also for general multi-level engines (see Appendix B for details).
+The quantity (3) is maximized when Alice knows the prior $P$ and sets $Q^{A}=P$, in which case $W=k_{B}TD(P||Q^{B})$.
+This maximal extractable work from the system can be identified with the nonequilibrium free energy from thermodynamics up to a change in the base of the logarithm.
+In the particular case where Bob puts the partition in the middle, one obtains $k_{B}T(\ln 2-H(P))$, where $H(\cdot)$ is the Shannon entropy.
+The expression for the average extracted work by Alice is formally related to a formula appearing in the classical Kelly betting problem from information theory Kelly (1956): both involve a difference of relative entropies, and are equivalent up to an exponentiation.
+We develop this connection in detail in a companion paper Arcos et al.
+(2025).
+However, we note that, in Kelly’s gambling problem, wealth evolves multiplicatively, making the natural utility functions those of the CRRA type, as considered in Bleuler et al.
+(2020).
 It is also worth noting that Eq.
-(1) has also appeared in the work of [25] in the context of work extraction with an incorrect prior distribution.
-When Alice does not know the correct prior, our adversarial Szilard engine reduces to the formulation of [25] at the level of the ensemble average.
-However, here we are interested in the regime of finite-size $n$ without ensemble averages, and in the case where Alice knows the prior distribution.
-This will allow us to connect Eq.
+(3) has also appeared in Ref.
+Kolchinsky and Wolpert (2017) in the context of work extraction with an incorrect prior distribution (see also Refs.
+Kolchinsky and Wolpert (2021a, b)).
+When Alice does not know the correct prior, our adversarial Szilard engine reduces to the formulation of Kolchinsky and Wolpert (2017) at the level of the ensemble average.
+We note that the ensemble average is reproduced from the empirical average in the limit of infinitely large sampling number $n\to\infty$ (assuming i.i.d.), while we will focus on the regime of finite sampling size $n$ in Section IV.
+Importantly, the primary focus of the subsequent arguments will be connecting Eq.
 (3) to expected utility theory, economics, and Kelly betting, all of which assume a known distribution in the ideal case.
-We will thus assume that the prior distribution is known to Alice.
+Therefore, we will assume that the prior distribution is known to Alice.
 Readers unfamiliar with Kelly betting may find it helpful to read the review in Appendix A before proceeding to the next section.
-III Utility maximisation: connection to expected utility formulations
+III Utility maximization: connection to expected utility formulations
 In the previous section, we argued that the Kelly gambling problem and the adversarial Szilard engine work extraction problem were mathematically equivalent up to an exponentiation.
@@
 In this sense, the optimal strategy for different agents depends on their particular level of risk aversion.
-A notable example connecting the Kelly utility function to other forms of risk aversion is given in [26], where the authors analyze Kelly gambling from the perspective of Constant Relative Risk Aversion (CRRA).
+A notable example connecting the Kelly utility function to other forms of risk aversion is given in Bleuler et al.
+(2020), where the authors analyze Kelly gambling from the perspective of Constant Relative Risk Aversion (CRRA).
 CRRA describes a class of utility functions in which an individual’s relative risk aversion remains constant regardless of their level of wealth.
@@
 This property makes CRRA particularly suitable for analyzing the Kelly paradigm, as it aligns with the proportional nature of Kelly betting, since it matches financial scenarios where wealth grows multiplicatively through compounding returns.
-On the other hand, in thermodynamics and Szilard engines, the system responds to changes in fixed energy amounts.
-Losing a fixed amount of energy matters the same regardless of total system energy- this is captured in the economics literature by the CARA utility function.
-This utility function has already proven useful in other thermodynamic contexts [16].
+On the other hand, in thermodynamics, the system responds to changes in fixed energy amounts.
+Losing a fixed amount of energy matters the same regardless of total system energy — this is captured in the economics literature by the CARA utility function.
+This utility function has already proven useful in other thermodynamic contexts Ducuara et al.
+(2023).
 The distinction between the two types of risk aversion reflects how thermodynamic systems respond to additive shocks, whilst financial systems experience multiplicative shocks (where gains and losses compound like interest).
 We now explain how varying risk aversion plays a role in thermodynamic protocols.
-For readers without an economics background, let us begin by recalling that in expected utility theory, the most fundamental postulate is (called the expected utility hypothesis) states that rational individuals make decisions in order to maximise their expected utility, rather than expected value [27].
-In this context, risk aversion is reflected in the utility function through the concavity - which reflects the fact that a risk-averse person might prefer a guaranteed $\mathdollar 100$ over a 50 per cent chance of winning $\mathdollar 250$, even though the latter has a higher expected value.
+For readers without an economics background, let us begin by recalling that in expected utility theory, the most fundamental postulate is (called the expected utility hypothesis) states that rational individuals make decisions in order to maximize their expected utility, rather than expected value Mas-Colell et al.
+(1995).
+In this context, risk aversion is reflected in the utility function through the concavity — which reflects the fact that a risk-averse person might prefer a guaranteed $\mathdollar 100$ over a 50 per cent chance of winning $\mathdollar 250$, even though the latter has a higher expected value.
 Risk-seeking tendencies are defined analogously.
@@
 The utility function we are considering in the thermodynamic context (CARA) is given by
-$u_{r}(w_{x})=\frac{1}{r}(1-\exp(-rw_{x}))$ (2)
-Here, the parameter $r$ reflects the risk preferences of the gambler
-  • $r>0$: Risk-aversion (concave utility) - the agent prefers guaranteed work extraction over uncertain fluctuations
-  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations
-  • $r<0$: Risk-seeking behaviour (convex utility) - agents prefer higher amount of work extracted even if it means lower probability of success
-As it is known from elementary thermodynamics [28], for a single round of work extraction in the generalised Szilard engine described above, the extracted work in terms of the outcome $x$ of the random variable $X$ can be expressed as
-$w_{x}=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (3)
-It follows using eq. 2 that the utility of outcome $x$ is
-$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (4)
-By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
-$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (5)
-In appendix D, we show that for a given $r$, choice of $Q_{X}^{A}$ corresponds to
-$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (6)
-Where we have introduced the notation $Q_{X}^{A,r}$ to indicate that the optimiser depends on the risk aversion of the gambler.
-Note that as $r$ goes to zero (risk neutral) $Q_{X}^{A,r}$ goes to $P_{X}$, and as $r$ goes to $\infty$ (extreme risk aversion) $Q_{X}^{A,r}$ goes to $Q_{X}^{B}$ (not moving partition, and hence no fluctuations at all), as expected.
-In the following section, we connect the strategy $Q_{X}^{A,r}$ derived above the average work extracted by a gambler with a particular level of risk aversion, as well as to Renyi divergences (the generalised free energies of the resource-theoretic approach).
-III.1 Risk aversion revisited
-The aim of this section is to revisit risk aversion in the thermodynamic setting and to show explicitly how the economic notion of risk aversion helps provide the missing bridge between stochastic and resource-theoretic formulations of the second law.
-Recall that, in the previous section, the utility maximisation problem of the adversarial Szilard engine led to optimal strategies for Alice of the form
-$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (7)
-In our engine, this corresponds directly to Alice’s choice of partition placement; in economic terms, it reflects the unique optimal allocation for a CARA agent with parameter $r$.
-The parameter $r$ controls the risk aversion of the gambler [29] as follows:
-  • $r>0$: Risk-aversion - the agent prefers guaranteed work extraction over uncertain fluctuations. Here risk aversion increases with increasing $r$.
-  • $r=0$: Risk-neutral behaviour (linear utility, recoverable via limit $r\rightarrow 0$) - agents maximize expected work regardless of fluctuations.
-  • $r<0$: Risk-seeking behaviour - agents prefer higher amount of work extracted even if it means lower probability of success.
-It is useful to evaluate the expected work extracted for a particular strategy.
-In appendix F, we show that for risk aversion level $r$, the expected work extracted $\mathbb{E}_{r}(Q_{X}^{A})$ is given by
-$\boxed{\mathbb{E}_{r}[W]=\big(\alpha D(P_{X}||Q_{X}^{B})+(1-\alpha)D_{\alpha}(P_{X}||Q_{X}^{B})\big)k_{B}T}$ (8)
-where we have included the subscript $\alpha=\frac{1}{1+r}$ to connect the level of risk aversion to the Renyi parameter and to indicate that the expected value of the extracted work depends on the particular strategy and level of risk aversion.
+$u_{r}(w)=\frac{1-e^{-rw/k_{B}T}}{r},$ (4)
+where $w$ is the work as a random variable. The parameter $r$ reflects the risk preferences of the gambler:
+  • $r>0$: Risk-aversion (concave utility) — The agent prefers guaranteed work extraction over uncertain fluctuations. Here risk aversion increases with increasing $r$.
+  • $r=0$: Risk-neutral behavior (linear utility, recoverable via limit $r\rightarrow 0$) — The agent maximizes expected work regardless of fluctuations.
+  • $r<0$: Risk-seeking behavior (convex utility) — The agent prefers higher amount of work extracted even if it means lower probability of success.
+For a single round of work extraction in the adversarial Szilard engine described above, the extracted work in terms of the outcome $x$ can be expressed as
+$w(x)=k_{B}T\ln\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg).$ (5)
+See also Appendix B for the case of multi-level engines. It follows using Eq. (4) that the utility of outcome $x$ is
+$u_{r}(w(x))=\frac{1}{r}\bigg(1-\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg).$ (6)
+By the expected utility hypothesis, a rational gambler will set their bet $Q^{A}$ as to maximize
+$\sum_{x}P(x)u_{r}(w(x))=\frac{1}{r}\bigg(1-\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg).$ (7)
+In Appendix C, we show that for a given $r$, the optimal choice of $Q^{A}$ corresponds to
+$Q^{A,r}=\frac{P(x)^{\frac{1}{1+r}}Q^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P(x^{\prime})^{\frac{1}{1+r}}Q^{B}(x^{\prime})^{\frac{r}{1+r}}},$ (8)
+where we have introduced the notation $Q^{A,r}$ to indicate that the optimizer depends on the risk aversion of the gambler.
+Note that as $r$ goes to zero (risk neutral) $Q^{A,r}$ goes to $P$, and as $r$ goes to $\infty$ (extreme risk aversion) $Q^{A,r}$ goes to $Q^{B}$ (not moving partition, and hence no fluctuations at all), as expected.
+In our adversarial Szilard engine, the strategy $Q^{A,r}$ corresponds directly to Alice’s choice of partition placement; in economic terms, it reflects the unique optimal allocation for a CARA agent with parameter $r$.
+In the following, we connect $Q^{A,r}$ to the average work extracted by a gambler with a particular level of risk aversion, as well as to the Rényi divergences (the generalized free energies of the resource-theoretic approach).
+It is useful to evaluate the expected work extracted for a particular strategy. In Appendix E, we show that for risk aversion level $r$, the average extracted work, written as $W^{(r)}$, is given by
+$\boxed{W^{(r)}=k_{B}T\left(\frac{1}{1+r}D(P||Q^{B})+\frac{r}{1+r}D_{\frac{1}{1+r}}(P||Q^{B})\right)}$ (9)
+Here, we defined the Rényi divergence for $\infty<\alpha<\infty$ following to Ref. van Erven and Harremoës (2014) :
+$D_{\alpha}(P||Q^{B})=\frac{1}{\alpha-1}\ln\sum_{x}P(x)^{\alpha}Q^{B}(x)^{1-\alpha}.$ (10)
+Eq.
+(9) clarifies that the ensemble average of the extracted work depends on the particular strategy and level of risk aversion.
+Noticing that $\frac{1}{1+r}+\frac{r}{1+r}=1$, Eq.
+(9) is a weighted sum of $D(P||Q^{B})$ and $D_{\frac{1}{1+r}}(P||Q^{B})$.
 […]
-$\boxed{W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T}$ (9)
-An analogous expression does not hold in the context of gambling, due to the fact that the utility function in that case is different.
-This result is complementary to the result of  [16] who quantified the dissipated fluctuating work $W_{CE}^{\mathrm{diss},\,r}$ in a thermal system which starts in equilibrium, and is driven out of it.
-This is the Crooks’ fluctuation relation setting.
-There, they find
-$\beta W_{\mathrm{CE}}^{\mathrm{diss},\,r}=D_{1+r}\!\big(P_{\mathrm{F}}(w)\,\|\,P_{\mathrm{R}}(-w)\big),$ (10)
+$\boxed{W_{CE}=k_{B}TD_{\frac{1}{1+r}}(P||Q^{B})}$ (11)
+We note that the certainty equivalent for $n$ rounds is exactly $n$ times the single-round certainty equivalent due to the additivity of work and independence of rounds.
+An analogous expression does not hold in the context of gambling, because the utility function in that case is different.
+A related but distinct approach was taken in Ref.
+Ducuara et al.
+(2023), which quantified the dissipated fluctuating work $W_{CE}^{\mathrm{diss},\,r}$ for a system starting in thermal equilibrium and driven out of it via a protocol satisfying the Crooks fluctuation theorem.
+They found
+$\beta W_{CE}^{\mathrm{diss},\,r}=D_{1+r}\!\big(P_{\mathrm{F}}(w)\,\|\,P_{\mathrm{R}}(-w)\big),$ (12)
 where $P_{F}(w)$ is the probability of the forward process and $P_{F}(-w)$ is the probability of the reverse process.
-Owing to the difference in initial starting states, it is unclear if these two expressions can be related, but it does allow and agent to decide between two different work extraction games in terms of expected utility.
+The two results, Eqs.
+(11) and (12), are derived from different physical premises and are not directly comparable.
+Our result (11) applies to a non-equilibrium work extraction game (the adversarial Szilard engine) that does not satisfy Crooks’ theorem, as the work values are determined by the agents’ choices independently of the referee’s distribution.
+In contrast, the result of Ref.
+Ducuara et al.
+(2023) is rooted in the fluctuation-theoretic framework of Crooks, which relates forward and reverse processes for a system initially at equilibrium.
+Despite these different settings, both works highlight how Rényi divergences naturally emerge when incorporating risk sensitivity into thermodynamic reasoning.
 It is useful to study the certainty equivalent and the expected work side by side.
-The expectation captures the stochastic-thermodynamic viewpoint of average entropy production, while the certainty equivalent aligns with the resource-theoretic notion of generalised free energies.
-Analysing their joint behaviour therefore clarifies how different levels of risk aversion interpolate between the two frameworks, and provides a rationality check on the agent’s strategy.
-From the explanations above, we see that in the Szilard engine picture, the strategy for which the agent places the partition proportional to probability corresponds to risk neutrality $r=0$ in the economic formalism [29].
+The expectation captures the stochastic-thermodynamic viewpoint of average entropy production, while the certainty equivalent aligns with the resource-theoretic notion of generalized free energies.
+Analyzing their joint behavior therefore clarifies how different levels of risk aversion interpolate between the two frameworks, and provides a rationality check on the agent’s strategy.
+From the explanations above, we see that in the adversarial Szilard engine picture, the strategy for which the agent places the partition proportional to probability corresponds to risk neutrality $r=0$ in the economic formalism Abbas (2018).
 Indeed, in expected utility theory, a risk neutral individual is one for which the certainty equivalent equals the expected value of their gamble.
-For this risk attitude, both correspond to the thermodynamic non-equilibrium free energy.
+For this risk attitude, both correspond to the thermodynamic nonequilibrium free energy.
 Let us now focus on strategies corresponding to $r\geq 0$, i.e.
 risk aversion.
-Here the certainty equivalent $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})$ eq.
-41 is non-negative, which follows directly from properties of the Renyi divergence in this regime [1].
-Note also that since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the certainty equivalent decreases with increasing risk aversion, as expected.
-Since the Rényi divergence $D_{\alpha}$ is a non-decreasing function of its order $\alpha$ [30], and since $\alpha<1$ for $r>0$, it follows that $D_{\alpha}(P_{X}\|Q_{X}^{B})\leq D_{1}(P_{X}\|Q_{X}^{B})=D(P_{X}\|Q_{X}^{B})$.
-The weight $(1-\alpha)$ is positive, ensuring the second term is non-negative.
-However, because $D_{\alpha}(P_{X}\|Q_{X}^{B})$ is strictly less than $D(P_{X}\|Q_{X}^{B})$ for non-trivial distributions ($P_{X}\neq Q_{X}^{B}$), the weighted sum $\mathbb{E}[W]$ is also strictly less than the maximum $D(P_{X}\|Q_{X}^{B})$.
-This shortfall compared to the non-equilibrium free energy is the price the agent pays for certainty.
-A subtlety arises for partition placements corresponding to risk-seeking behaviour ($r<0$), where the sign of both the certainty equivalent and the expected value may shift.
+Here the certainty equivalent $D_{\frac{1}{1+r}}(P||Q^{B})$ [Eq.
+(11)] is non-negative, which follows directly from properties of the Rényi divergence in this regime Sagawa (2022).
+Note also that since the Rényi divergences satisfy van Erven and Harremoës (2014)
+$D_{\alpha_{1}}(P||Q^{B})\leq D_{\alpha_{2}}(P||Q^{B})\ \ \ \text{for any}\ -\infty\leq\alpha_{1}<\alpha_{2}\leq\infty,$ (13)
+the certainty equivalent decreases with increasing risk aversion, as expected.
+Also, since $\frac{1}{1+r}<1$ for $r>0$, it follows that $D_{\frac{1}{1+r}}(P\|Q^{B})\leq D_{1}(P\|Q^{B})=D(P\|Q^{B})$.
+The weight $\frac{r}{1+r}$ on the right-hand side of Eq.
+(9) is positive, ensuring the second term is non-negative.
+However, because $D_{\frac{1}{1+r}}(P\|Q^{B})$ is strictly less than $D(P\|Q^{B})$ for nontrivial distributions ($P\neq Q^{B}$), the weighted sum $W^{(r)}$ is also strictly less than the maximum $D(P\|Q^{B})$.
+This shortfall compared to the nonequilibrium free energy is the price the agent pays for certainty.
+A subtlety arises for partition placements corresponding to risk-seeking behavior ($r<0$), where the sign of both the certainty equivalent and the expected value may shift.
 Thermodynamically, risk-seeking strategies with negative expected value mark an operating point at which the agent aims to violate the second law: while rare fluctuations might yield temporary gains, they fail on average.
-The rationality of the agent in such regimes is described by the joint behaviour of the certainty equivalent and the expected value in the following sense: if the certainty equivalent is more negative than the expected value of the gamble, the preferences remain consistent with expected utility theory.
+The rationality of the agent in such regimes is described by the joint behavior of the certainty equivalent and the expected value in the following sense: if the certainty equivalent is more negative than the expected value of the gamble, the preferences remain consistent with expected utility theory.
 In this case, the agent is simply expressing a preference for the risky gamble which has a non-zero probability of a gain over a certain loss.
-This is consistent with risk-seeking behaviour.
-However, risk-seeking behaviour becomes pathological if the certainty equivalent is more negative than the gamble’s worst possible outcome.
+This is consistent with risk-seeking behavior.
+However, risk-seeking behavior becomes pathological if the certainty equivalent is more negative than the gamble’s worst possible outcome.
 In this scenario, the guaranteed loss prescribed by the certainty equivalent is strictly worse than every possible result of the gamble itself.
-This constitutes a violation of first-order stochastic dominance—a fundamental axiom of rational choice—as the dominated sure loss should never be chosen over the gamble that uniformly outperforms it [27, 31].
-Hence, for risk-seeking behaviour, we must analyse the cases 1<r<0-1<r<0 and r<1r<-1 separately.
-For 1<r<0-1<r<0, the certainty equivalent remains positive and greater than the non-equilibrium free energy, reflecting the fact that the agent would only give up the gamble for a guaranteed amount greater than the non-equilibrium free energy.
+This constitutes a violation of first-order stochastic dominance — a fundamental axiom of rational choice — as the dominated sure loss should never be chosen over the gamble that uniformly outperforms it Mas-Colell et al.
+(1995); Ingersoll (1987).
+That is, the violation of the first-order stochastic dominance occurs if the certainty equivalent is more negative than the worst-case work output:
+$W_{CE}<\min_{x}w(x).$ (14)
+Hence, for risk-seeking behavior, we must analyze the cases 1<r<0-1<r<0 and r<1r<-1 separately.
+For 1<r<0-1<r<0, the certainty equivalent remains positive and greater than the nonequilibrium free energy, reflecting the fact that the agent would only give up the gamble for a guaranteed amount greater than the nonequilibrium free energy.
 For r<1r<-1, where the certainty equivalent flips sign, we must verify the stochastic dominance condition.
-In appendix G, we show that in our generalised Szilard engine, the condition for a violation of stochastic dominance in the regime where r<1r<-1 (the regime for which the sign of the certainty equivalent flips) is given by
-$D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$ (11)
-However, the right hand side is equal to $D_{-\infty}(P_{X}||Q_{X}^{B})$, so the condition reduces to $D_{\frac{1}{1+r}}(P_{X}||Q_{X}^{B})<D_{-\infty}(P_{X}||Q_{X}^{B})$.
-But since for all $\alpha\in(-\infty,\infty)$ the Renyi divergences satisfy $D_{\alpha_{1}}(P_{X}||Q_{X}^{B})\leq D_{\alpha_{2}}(P_{X}||Q_{X}^{B})$ for $\alpha_{1}<\alpha_{2}$, the first-order stochastic dominance condition is never violated.
+In Appendix F, we show that in our adversarial Szilard engine, the condition for a violation of stochastic dominance in the regime where r<1r<-1 is given by
+$D_{\frac{1}{1+r}}(P||Q^{B})>\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right).$ (15)
+Since the right-hand side is equal to $D_{\infty}(P||Q^{B})$, the condition reduces to $D_{\frac{1}{1+r}}(P||Q^{B})>D_{\infty}(P||Q^{B})$.
+But from inequality (13), this never happens, and thus the first-order stochastic dominance condition is never violated.
 IV Thermodynamic risk aversion
-In the section II, we showed that the non-equilibrium free energy from thermodynamics can be understood as the maximum amount of work that can be extracted in the presence of an adversary at the level of the ensemble average.
-The aim of this section is to analyse the same ensemble from a finite-size regime perspective using tools developed in [17], and to prove that work extraction in the finite-size regime corresponds to the maximisation of a utility function as described in section III.
-Recall first that for a single round of work extraction, the extracted work in terms of the outcome $x$ of the random variable can be expressed as
-$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T.$ (12)
-Again, this formula is true not only for the above binary adversarial Szilard engine (Fig. 1), but also for general multi-level engines (see Appendix B for details).
-It follows that for $n$ rounds, the work extracted is given by
-$W_{n}=\sum_{i=1}^{n}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (13)
-by grouping terms with the same outcome, and denoting by $N_{x}$ the number of times each outcome occurs, this can be expressed as
-$W_{n}=\sum_{x}\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}k_{B}T$ (14)
-It is useful to define the type $\lambda_{x^{n}}$ of the sequence $x^{n}$.
-Often called its empirical distribution, the type is defined as $\lambda_{x^{n}}(x)=N_{x}/n$.
+In Section II, we showed that the nonequilibrium free energy from thermodynamics can be understood as the maximum amount of work that can be extracted in the presence of an adversary at the level of the ensemble average.
+The aim of this section is to analyze the same ensemble from a finite-size regime perspective using tools developed in Arcos et al.
+(2025), and to prove that work extraction in the finite-size regime corresponds to the maximization of a utility function as described in Section III.
+Recall first that for a single round of work extraction, the extracted work in terms of the outcome $x$ of the random variable is expressed by Eq.
+(5).
+It follows that for $n$ rounds of i.i.d., the work extracted is given by
+$W_{n}(x^{n})=k_{B}T\sum_{i=1}^{n}\ln\bigg(\frac{Q^{A}(x_{i})}{Q^{B}(x_{i})}\bigg),$ (16)
+where $x_{i}$ is the outcome of the $i$th round and $x^{n}=(x_{1},x_{2},\cdots,x_{n})$ describes the realized sequence.
+By grouping terms with the same outcome, and denoting by $N_{x^{n}}(x)$ the number of times each outcome occurs in $x^{n}$, this can be expressed as
+$W_{n}(x^{n})=k_{B}Tn\sum_{x}N_{x^{n}}(x)\ln\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg).$ (17)
+It is useful to introduce the type $\lambda_{x^{n}}$ of the sequence $x^{n}$ Csiszár (1998).
+Often called its empirical distribution, the type is defined as $\lambda_{x^{n}}(x)=N_{x^{n}}(x)/n$.
 Rewriting the equation above in terms of the type, we obtain an equation analogous to Eq.
-1 in the finite-size regime:
-$\boxed{W_{n}=n(D(\lambda_{x^{n}}||Q^{B}_{X})-D(\lambda_{x^{n}}||Q^{A}_{X}))k_{B}T}$ (15)
-Hence, by placing the partition at position $Q_{X}^{A}=\lambda_{x^{n}}$ (when Alice’s allocation matches the type of the observed sequence), the work extracted is given by $n(D(\lambda_{x^{n}}||Q^{B}_{X}))k_{B}T$.
-In the finite-size regime, the work extraction problem reduces to what is essentially a decision-theoretic problem of “guessing the type” of the empirical sequence $\lambda_{x^{n}}$ that will be realised.
-In concurrent work [17], we show that each possible type corresponds to a particular fluctuation, with probability decaying exponentially up to sub-exponential factors as $P_{X}(\lambda_{x^{n}})\doteq\exp(-nD(\lambda_{x^{n}}\|P_{X}))$, while the potential work reward scales as $nD(\lambda_{x^{n}}\|Q_{X}^{B})$.
-This creates a fundamental risk-reward trade-off: Alice must select her strategy $Q_{X}^{A}$ to balance the probability of success against the amount of work extracted when successful.
-In appendix C, we show that the optimal strategy that maximises work extraction for a given risk tolerance (minimum success probability $\epsilon$) takes the form of an exponential family interpolating between $P_{X}$ and $Q_{X}^{B}$:
-$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (16)
-where the parameter $\mu$ is determined by the risk constraint. The corresponding work bound is given in terms of a Rényi divergence:
-$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (17)
-The detailed derivation using the method of types [32] is based on a treatment of Kelly gambling in the finite-size regime developed in concurrent work [17].
-This approach reveals that all Rényi divergences $D_{\alpha}$ acquire an operational interpretation in terms of work extraction for different risk tolerances.
-Previous work [20] provided operational interpretation for min and max entropies as bounds in finite-size work extraction.
-Here, we find similar operational interpretation for all Renyi divergences.
+(3) in the finite-size regime:
+$\boxed{W_{n}(x^{n})=k_{B}Tn(D(\lambda_{x^{n}}||Q^{B})-D(\lambda_{x^{n}}||Q^{A}))}$ (18)
+Hence, by placing the partition at position $Q^{A}=\lambda_{x^{n}}$ (when Alice’s allocation matches the type of the observed sequence), the work extracted is given by $k_{B}TnD(\lambda_{x^{n}}||Q^{B})$.
+In the finite-size regime, the work extraction problem reduces to what is essentially a decision-theoretic problem of “guessing the type” of the empirical sequence $\lambda_{x^{n}}$ that will be realized.
+Each possible type corresponds to a particular fluctuation with probability decaying exponentially up to sub-exponential factors, while the potential work reward scales as $nD(\lambda_{x^{n}}\|Q^{B})$.
+More precisely, the probability of a type is bounded as Cover and Thomas (2006)
+$P(\lambda_{x^{n}})\leq\exp(-nD(\lambda_{x^{n}}\|P)).$ (19)
+This creates a fundamental risk-reward trade-off: Alice must select her strategy to balance the probability of success against the amount of work extracted when successful.
+Concretely, Alice’s strategy is to match her choice $Q^{A}$ to a particular realization of type $\lambda_{x^{n}}$ with success probability larger than a given threshold $\epsilon>0$, that is,
+$P(\lambda_{x^{n}})\geq\epsilon,$ (20)
+which determines the risk tolerance. From this and Eq. (19), the risk constraint can be represented as
+$D(\lambda_{x^{n}}\|P)\leq\frac{1}{n}\ln\left(\frac{1}{\epsilon}\right).$ (21)
+Under this constraint, we can derive the optimal strategy that maximizes work, $W_{n}(x^{n})=k_{B}TnD(\lambda_{x^{n}}||Q^{B})$.
+Then, the optimization problem is phrased as: Maximize the reward $D(\lambda_{x^{n}}\|Q^{B})$ with respect to $\lambda_{x^{n}}$ ($Q^{A}$) subject to the constraint (21).
+Note that Alice does not know which type $\lambda_{x^{n}}$ is actually realized beforehand; she bets on $Q^{A}$ and will be successful if $\lambda_{x^{n}}=Q^{A}$ is realized.
+Using the standard Lagrange multiplier method, we obtain the family of optimal strategies, taking the form of an exponential family interpolating between $P$ and $Q^{B}$:
+$Q^{A*,\mu}(x)=\frac{P(x)^{\mu}Q^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P(x^{\prime})^{\mu}Q^{B}(x^{\prime})^{1-\mu}},$ (22)
+where the parameter $0\leq\mu\leq 1$ is determined by the risk constraint $\epsilon$.
+See also Section 11.7 of Ref.
+Cover and Thomas (2006) for a formally similar optimization problem related to hypothesis testing.
+The corresponding work bound is given in terms of the Rényi divergence:
+$W_{n}(x^{n})\geq k_{B}T\left(nD_{\mu}(P\|Q^{B})+\frac{\mu}{1-\mu}\ln\epsilon\right).$ (23)
+While this is also argued in the concurrent work Arcos et al.
+(2025), we show its derivation in Appendix G for self-containedness.
+We note that there does not necessarily exist a type $\lambda_{x^{n}}$ that exactly matches $Q^{A*,\mu}$ defined by Eq.
+(22) for finite $n$ (especially when $Q^{A*,\mu}(x)$ is an irrational number), and thus the above argument should involve some approximation for finite $n$.
+Our approach reveals that all Rényi divergences $D_{\alpha}$ acquire an operational interpretation in terms of work extraction for different risk tolerances.
+Previous study Dahlsten et al.
+(2011) provided operational interpretation for min and max entropies as bounds in finite-size work extraction.
+Here, we find another operational interpretation for all Rényi divergences.
 Crucially, we also provide the work extraction strategy given the agent’s constraints.
-In addition, this result is identical to the expected utility formulation of the problem after identifying
-$\frac{r}{1+r}=1-\frac{1}{1+r}$ (18)
-with $\mu$ in eq. 16.
+In addition, this result is identical to the expected utility formulation (8), by identifying $\mu$ to $\frac{1}{1+r}$.
 V Conclusions
@@
 In fact, the certainty equivalents defined by CARA utilities coincide exactly with Rényi divergences, demonstrating that the two modern perspectives — the stochastic description of fluctuating work and the resource-theoretic hierarchy of free energies — are not separate constraints but two ways of expressing the same decision-theoretic principles.
-While the generalised free energies found in the context of second laws are given in terms of the Renyi-divergence from the initial state $P_{X}$ to the equilibrium state at inverse temperature $\beta$[8], here we found that the relevant Renyi-divergences is from the state $P_{X}$ to the out of equilibrium state $Q_{X}^{B}$.
-Understanding the relationships between these, is an interesting open question.
-Acknowledgements
+Acknowledgments
 MA thanks Ian Ford and Paul Skrzypczyk for useful discussions.
@@
 PhF is supported by the project FOR 2724 of the Deutsche Forschungsgemeinschaft (DFG).
-TS is supported by by JST ERATO Grant Number JPMJER2302, Japan.
+TS is supported by JST ERATO Grant Number JPMJER2302, Japan.
 JO was supported by the Simons Foundation, It from Qubit Network, and EPSRC.
 Appendix A Review of Kelly betting
-The Kelly [18] betting scheme is an adversarial set-up in which the gambler, Alice, allocates fractions of her wealth to the possible outcomes of a horse race, always distributing a non-zero fraction to each outcome (in order to avoid ever being completely broke).
-The adversary Bob, also sets odds for each possible outcome.
-Suppose that the horse race is described by the random variable $X$.
+The Kelly Kelly (1956) betting scheme is an adversarial setup in which the gambler, Alice, allocates fractions of her wealth to the possible outcomes of a horse race, always distributing a non-zero fraction to each outcome (in order to avoid ever being completely broke).
+The adversary Bob also sets odds for each possible outcome.
+Suppose that the horse race is described by a random variable $\{x\}$.
 Bob, sets the odds $o_{x}$ to outcome $x$, and Alice allocates the fraction $f_{x}$ of her wealth to outcome $x$.
-With this convention for describing odds, after a round of the game, Alice’s initial wealth $W_{i}$ is multiplied by the factor $f_{x}o_{x}$:
-$W_{1}=f_{x}o_{x}W_{i}$ (19)
-Assuming that Alice re-invests whatever wealth she has after the previous round, her wealth evolves recursively, and at round $N$ it obeys the equation
-$W_{n}=f_{x}o_{x}W_{n-1}$ (20)
+With this convention for describing odds, after a round of the game, Alice’s initial wealth $\mathcal{W}_{i}$ is multiplied by the factor $f_{x}o_{x}$:
+$\mathcal{W}_{1}=f_{x}o_{x}\mathcal{W}_{i}.$ (24)
+Assuming that Alice reinvests whatever wealth she has after the previous round, her wealth evolves recursively, and at round $N$ it obeys the equation
+$\mathcal{W}_{n}=f_{x}o_{x}\mathcal{W}_{n-1}.$ (25)
 Since Alice reinvests her wealth after each round, her wealth after $n$ rounds can be expressed as a product of the wealth multipliers for each outcome:
-$\frac{W_{n}}{W_{i}}=\prod_{x}(f_{x}o_{x})^{N_{x}}$ (21)
-where $N_{x}$ is the number of times that the outcome $x$ occurred.
-Since the fraction by which the wealth is multiplied after every round is a random variable, it is customary to write $Q_{X}^{B}(x)=o_{x}^{-1}$ and $Q^{A}_{X}(x)=f_{x}$.
-Here, the subscript is there to $X$ indicate that the ratio $\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}$ is a random variable, whilst the superscripts $A$ and $B$ indicate that the fraction is chosen by Alice (resp.
-Bob).
-Altogether, the notation helps to stress the fact that even though the bets and odds are the same at every round of gambling, the amount of money made by Alice is a random variable.
+$\frac{\mathcal{W}_{n}}{\mathcal{W}_{i}}=\prod_{x}(f_{x}o_{x})^{N_{x^{n}}(x)},$ (26)
+where $N_{x^{n}}(x)$ is the number of times that the outcome $x$ occurred in the realized sequence $x^{n}=(x_{1},x_{2},\cdots,x_{n})$.
+Since the fraction by which the wealth is multiplied after every round is a random variable, it is customary to write $Q^{B}(x)=o_{x}^{-1}$ and $Q^{A}(x)=f_{x}$.
+The superscripts $A$ and $B$ indicate that the fraction is chosen by Alice and Bob, respectively.
+Even though the bets and odds are the same at every round of gambling, the amount of money made by Alice is a random variable.
 The expression above now becomes
-$W_{n}=\prod_{x}\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{N_{x}}W_{i}$ (22)
-In the limit as $n>>1$, the ratio of Alice’s initial wealth $W_{i}$ to her final wealth $W_{F}$ satisfies:
-$\frac{W_{F}}{W_{i}}=\exp(n(D(P_{X}||Q_{X}^{B})-D(P_{X}||Q_{X}^{A})))$ (23)
-Here, $D(P_{X}\|Q_{X})$ is the relative entropy (or Kullback-Leibler divergence).
-Though not a true distance metric, it is often interpreted in information theory as measuring the discrepancy between probability distributions.
-Since relative entropy is non-negative, Alice’s optimal strategy is to set $Q_{X}^{A}=P_{X}$, ensuring that her wealth grows at the maximum possible rate.
-The expression above is the well-known result by Kelly [18].
+$\mathcal{W}_{n}=\prod_{x}\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{N_{x^{n}}(x)}\mathcal{W}_{i}.$ (27)
+In the limit as $n\gg 1$, the ratio of Alice’s initial wealth $W_{i}$ to her final wealth $\mathcal{W}_{F}$ satisfies
+$\frac{\mathcal{W}_{F}}{\mathcal{W}_{i}}=\exp(n(D(P||Q^{B})-D(P||Q^{A}))).$ (28)
+Since the relative entropy is non-negative, Alice’s optimal strategy is to set $Q^{A}=P$, ensuring that her wealth grows at the maximum possible rate.
+The expression above is the well-known result by Kelly Kelly (1956).
 This multiplicative growth of wealth under Kelly gambling contrasts with the additive accumulation of work in the thermodynamic engine, fundamentally shaping the respective utility functions (CRRA vs.
@@
 Appendix B General protocol for adversarial work extraction
-In this section, we derive the work formula for a general adversarial engine.
+In this appendix, we derive the work formula for a general adversarial engine.
 Let $x$ be a microscopic state that is not necessarily binary.
-Suppose that the initial distribution is given by $P_{X}(x)$ and the initial energy level is given by $E_{X}^{B}(x)$.
-We relate the energy level to Bob’s distribution by $Q_{X}^{B}(x)=e^{-E_{X}^{B}(x)}/Z_{X}^{B}$ with $Z_{X}^{B}=\sum_{x}e^{-E_{X}^{B}(x)}$ being the normalization factor (partition function).
+Suppose that the initial distribution is given by $P(x)$ and the initial energy level is given by $E^{B}(x)$.
+We relate the energy level to Bob’s distribution by $Q^{B}(x)=e^{-E^{B}(x)}/Z^{B}$ with $Z^{B}=\sum_{x}e^{-E^{B}(x)}$ being the normalization factor (partition function).
 Note that we set $k_{B}T$ to unity throughout this section.
-Let us first remember the optimal work extraction protocol at the level of ensemble average: (i) Alice quenches the energy level to $E_{X}(x)$ that is defined through $P_{X}(x)=e^{-E_{X}(x)}/Z_{X}$ with $Z_{X}=\sum_{x}e^{-E_{X}(x)}$; (ii) Alice moves the energy level from $E_{X}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally.
-It is well known that Alice extracts the work $D(P_{X}\|Q_{X}^{B})k_{B}T$ on average by this protocol [1, 3].
+Let us first remember the optimal work extraction protocol at the level of ensemble average: (i) Alice quenches the energy level to $E(x)$ that is defined through $P(x)=e^{-E(x)}/Z$ with $Z=\sum_{x}e^{-E(x)}$; (ii) Alice moves the energy level from $E(x)$ to $E^{B}(x)$ quasi-statically and isothermally.
+It is well known that Alice extracts the work $D(P\|Q^{B})$ on average by this protocol Sagawa (2022); Parrondo et al.
+(2015).
 Now we turn to the adversarial scenario.
-The protocol that Alice performs is: (i) Alice quenches the energy level to $E_{X}^{A}(x)$ that is defined through $Q_{X}^{A}(x)=e^{-E_{X}^{A}(x)}/Z_{X}^{A}$ with $Z_{X}^{A}=\sum_{x}e^{-E_{X}^{A}(x)}$ for her choice of $Q_{X}^{A}$; (i’) Alice lets the system thermalized; (ii) Alice moves the energy level from $E_{X}^{A}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally.
-The extracted work in step (i) is given by $E_{X}^{B}(x)-E_{X}^{A}(x)$, while in step (ii) $\ln Z_{X}^{B}-\ln Z_{X}^{A}$ (the change in the equilibrium free energy).
+The protocol that Alice performs is: (i) Alice quenches the energy level to $E^{A}(x)$ that is defined through $Q^{A}(x)=e^{-E^{A}(x)}/Z^{A}$ with $Z^{A}=\sum_{x}e^{-E^{A}(x)}$ for her choice of $Q^{A}$; (i’) Alice lets the system thermalize; (ii) Alice moves the energy level from $E^{A}(x)$ to $E^{B}(x)$ quasi-statically and isothermally.
+The extracted work in step (i) is given by $E^{B}(x)-E^{A}(x)$, while in step (ii) $\ln Z^{B}-\ln Z^{A}$ (the change in the equilibrium free energy).
 We thus obtain
-$w_{X}(x)=E_{X}^{B}(x)-E_{X}^{A}(x)+\ln Z_{X}^{B}-\ln Z_{X}^{A}=\ln\left(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\right),$ (24)
-which reproduces Eq. (12). By taking the average with respect to $P_{X}$, Eq. (1) is also reproduced.
-Appendix C Work extraction and the method of types
-This appendix summarizes the application of the method of types to the adversarial Szilard engine, following the general approach detailed in concurrent work [17].
-The method of types [32] provides a powerful framework for analyzing finite-size fluctuations in information-theoretic tasks.
-For a sequence of $n$ rounds, the empirical distribution (type) $\lambda_{x^{n}}$ occurs with probability approximately $\exp(-nD(\lambda_{x^{n}}\|P_{X}))$.
-The work extracted when the type is $\lambda_{x^{n}}$ is $nD(\lambda_{x^{n}}\|Q_{X}^{B})k_{B}T$ when Alice sets $Q_{X}^{A}=\lambda_{x^{n}}$.
-The optimization problem becomes: maximize the reward $D(\lambda_{x^{n}}\|Q_{X}^{B})$ subject to the constraint that the probability of success exceeds $\epsilon$, i.e., $D(\lambda_{x^{n}}\|P_{X})\leq\frac{1}{n}\ln(1/\epsilon)$.
-Using Lagrange multipliers, one obtains the family of optimal strategies:
-$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (25)
-where $\mu\in[0,1]$ is determined by the constraint. Substituting back gives the work bound:
-$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (26)
-For a comprehensive treatment of the method of types in adversarial scenarios, including the connection to hypothesis testing and utility theory, we refer to [17, 21].
-Appendix D Maximisation of the utility function
-In this section, we compute the maximiser of the utility function for risk parameter $r$.
-Recall now that the extracted work from the adversarial Szilard engine for a single round can be described by the random variable
-$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (27)
-It follows using eq. 2 that the utility of outcome $x$ is
-$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (28)
-By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
-$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (29)
-Note that the first term is a constant, so we need only focus on the second term.
+$w(x)=E^{B}(x)-E^{A}(x)+\ln Z^{B}-\ln Z^{A}=\ln\left(\frac{Q^{A}(x)}{Q^{B}(x)}\right),$ (29)
+which reproduces Eq. (5). By taking the average with respect to $P$, Eq. (3) is also reproduced.
+Appendix C maximization of the utility function
+In this appendix, we compute the maximizer of the utility function for risk parameter $r$.
+Recall that a rational gambler will set their bet $Q^{A}$ as to maximize Eq.
+(7).
+Note that the first term on the right-hand side is a constant, so we need only focus on the second term.
 For $r>0$ (risk-aversion), the second term is negative, whilst if $r<0$ (risk-seeking), the second term is positive.
-This means that for a risk-averse individual, we minimise
-$\min_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (30)
-which is equivalent to minimising
-$\min_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (31)
-whilst for a risk-seeking individual, we maximise
-$\max_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (32)
-which is equivalent to maximising
-$\max_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (33)
-In order to find the minimiser (resp. maximiser), we will use a lemma proved in [26], which states that
-$\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})=-r\big(D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})\big)$ (34)
+This means that for a risk-averse individual, we minimize
+$\min_{Q^{A}}\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r},$ (30)
+which is equivalent to minimizing
+$\min_{Q^{A}}\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}).$ (31)
+On the other hand, for a risk-seeking individual, we maximize
+$\max_{Q^{A}}\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r},$ (32)
+which is equivalent to maximizing
+$\max_{Q^{A}}\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}).$ (33)
+In order to find the minimizer (resp. maximizer), we will use a lemma proved in Bleuler et al. (2020), which states that
+$\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r})=-r\big(D_{\frac{1}{1+r}}(P||Q^{B})-D_{1+r}(G^{r}||Q^{A})\big),$ (34)
 where
-$G_{X}^{r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (35)
-The maximiser is therefore
-$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (36)
-Appendix E Calculation of the certainty equivalent
-At the beginning of this section, we said that two equivalent ways of characterising risk preferences were through the concavity/convexity of the utility function and the certainty equivalent.
-As shown in [16], the certainty equivalent has a clear information-theoretic interpretation in a thermodynamic context.
+$G^{r}=\frac{P(x)^{\frac{1}{1+r}}Q^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P(x^{\prime})^{\frac{1}{1+r}}Q^{B}(x^{\prime})^{\frac{r}{1+r}}}.$ (35)
+The minimizer (resp. maximizer) is therefore given by Eq. (8).
+Appendix D Calculation of the certainty equivalent
+Recall that two equivalent ways of characterizing risk preferences were through the concavity/convexity of the utility function and the certainty equivalent.
+As shown in Ducuara et al.
+(2023), the certainty equivalent has a clear information-theoretic interpretation in a thermodynamic context.
 We will also find it useful here to calculate the certainty equivalent of work for each value of the risk aversion parameter $r$ below.
-Recall that, for an agent exhibiting risk aversion, the certainty equivalent is the guaranteed amount of money that the individual would accept instead of taking a risky bet with different payoff.
-We then say that an individual with utility function $u(.)$ is risk averse if and only if their certainty equivalent is lower than the expected value of the risky lottery.
-The mathematical definition of the certainty equivalent is the number $W_{CE}\in\mathbb{R}$ such that $u_{r}(W_{CE})=E(u_{r}(w_{X}(x)))$, where $E$ denotes expected value.
-Since the utility of outcome $x$ is
-$u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (37)
-we must calculate the number $W_{CE}$ $u_{r}(W_{CE})=E(u_{r}(x))$, where $E$ denotes expected value, we therefore calculate
-$W_{CE}=u_{r}^{-1}\big(E(u_{r}(w))\big)\\ =u_{r}^{-1}\bigg(E\bigg(\frac{1}{r}(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =u_{r}^{-1}\bigg(\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (38)
-where in the last line we have used that the inverse function of $u_{r}(x)$ is $u_{r}^{-1}(x)=-\frac{1}{r}\ln(1-rx)$. Since
-$W_{CE}=-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (39)
-which equals
-$D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})$ (40)
-by the result of [26], we finally conclude that
-$W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T$ (41)
-when the expression for $Q_{X}^{A,r}$ is substituted into the expression for the certainty equivalent.
+Since the utility of outcome $x$ is given by Eq. (6), we calculate
+$W_{CE}$ $=u_{r}^{-1}\big(\mathbb{E}(u_{r}(w(x)))\big)$ (36)
+$=u_{r}^{-1}\bigg(\mathbb{E}\bigg(\frac{1}{r}(1-\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg)\bigg)$ (37)
+$=u_{r}^{-1}\bigg(\frac{1}{r}\bigg(1-\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg)\bigg)$ (38)
+$=-\frac{1}{r}\ln\bigg(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg),$ (39)
+where in the last line we have used that the inverse function of $u_{r}(x)$ is $u_{r}^{-1}(x)=-\frac{1}{r}\ln(1-rx)$. Again from the result of Bleuler et al. (2020), we obtain
+$W_{CE}=D_{\frac{1}{1+r}}(P||Q^{B})-D_{1+r}(G^{r}||Q^{A}).$ (40)
+We finally conclude Eq. (11), when the expression for $Q^{A,r}$ is substituted into the expression for the certainty equivalent.
 Since $D_{\alpha}<D_{\beta}$ for $\alpha<\beta$, one sees that for risk-averse (resp.
@@
 decreasing) risk aversion, as expected.
-In addition, the work certainty equivalent is smaller than $D_{1}$ (the non-equilibrium free energy) for a risk averse player, and equal to $D_{1}$ for a risk neutral player.
-In particular, for an extremely risk averse player ($r\to\infty$), the certainty equivalent tends to $D_{0}(P_{X}||Q_{X}^{B})$, which becomes zero when both distributions have the same support.
-Appendix F Extracted work as a function of strategy
-In section II, we showed that for a given strategy of Alice $Q_{X}^{A}$, the average work extracted in the adversarial Szilard engine is given by
-$\mathbb{E}[W]=\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T$ (42)
-Our goal is to express this for the optimal strategy $Q_{X}^{A}$ which from now on we write as $Q^{A,r}_{X}$, to emphasise dependence on the risk parameter of the CARA utility function.
+In addition, the work certainty equivalent is smaller than $D_{1}$ (the nonequilibrium free energy) for a risk averse player, and equal to $D_{1}$ for a risk neutral player.
+In particular, for an extremely risk averse player ($r\to\infty$), the certainty equivalent tends to $D_{0}(P||Q^{B})$, which becomes zero when both distributions have the same support.
+Appendix E Extracted work as a function of strategy
+Our goal of this Appendix is to express Eq.
+(3) for the optimal strategy $Q^{A}$ which from now on we write as $Q^{A,r}$, to emphasize dependence on the risk parameter of the CARA utility function.
 We substitute the explicit form of the optimal strategy into the second term.
-The optimal strategy for a given risk parameter $r$ is:
-$Q^{A,r}X(x)=\frac{P_{X}(x)^{\frac{1}{1+r}}Q^{B}_{X}(x)^{\frac{r}{1+r}}}{Z},$ (43)
+Let $\alpha=\frac{1}{1+r}$ for notational clarity. The optimal strategy for a given risk parameter $r$ is given by
+$Q^{A,r}(x)=\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)},$ (41)
 where
-$Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q^{B}_{X}(x^{\prime})^{\frac{r}{1+r}}$ (44)
-Let $\alpha=\frac{1}{1+r}$ for notational clarity. Since the first divergence is independent of the strategy $Q^{A,r}_{X}$, we compute the second divergence term, $D(P_{X}||Q^{A,r}_{X})$:
-$D(P_{X}||Q^{A,r}_{X})$ $=\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{A,r}_{X}(x)}$ (45)
-$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\ln\left(\frac{P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}}{Z}\right)\right]$ (46)
-$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\alpha\ln P_{X}(x)-(1-\alpha)\ln Q^{B}_{X}(x)+\ln Z\right]$ (47)
-$=(1-\alpha)\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{B}_{X}(x)}+\ln Z$ (48)
-$=(1-\alpha)D(P_{X}||Q^{B}_{X})+\ln Z(\alpha)$ (49)
-The normalization constant $Z$ is related to the sum that defines the Rényi divergence. Recall its definition:
-$D_{\alpha}(P_{X}||Q^{B}_{X})=\frac{1}{\alpha-1}\ln\sum_{x}P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}=\frac{1}{\alpha-1}\ln Z$ (50)
-So that $\ln Z$:
-$\ln Z(\alpha)=(\alpha-1)D_{\alpha}(P_{X}||Q^{B}_{X})$ (51)
+$Z(\alpha)=\sum_{x^{\prime}}P(x^{\prime})^{\alpha}Q^{B}(x^{\prime})^{1-\alpha}.$ (42)
+Since the first divergence is independent of the strategy $Q^{A,r}$, we compute the second divergence term:
+$D(P||Q^{A,r})$ $=\sum_{x}P(x)\ln\frac{P(x)}{Q^{A,r}(x)}$ (43)
+$=\sum_{x}P(x)\left[\ln P(x)-\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)}\right)\right]$ (44)
+$=\sum_{x}P(x)\left[\ln P(x)-\alpha\ln P(x)-(1-\alpha)\ln Q^{B}(x)+\ln Z(\alpha)\right]$ (45)
+$=(1-\alpha)\sum_{x}P(x)\ln\frac{P(x)}{Q^{B}(x)}+\ln Z(\alpha)$ (46)
+$=(1-\alpha)D(P||Q^{B})+\ln Z(\alpha).$ (47)
+From Eq. (10), the normalization constant $Z(\alpha)$ is related to the Rényi divergence:
+$\ln Z(\alpha)=(\alpha-1)D_{\alpha}(P||Q^{B}).$ (48)
 Substituting (48) back into (47):
-$D(P_{X}||Q^{A,r}X)$ $=(1-\alpha)D(P_{X}||Q^{B}_{X})+(\alpha-1)D\alpha(P_{X}||Q^{B}_{X})$ (52)
-$=(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]$ (53)
+$D(P||Q^{A,r})$ $=(1-\alpha)D(P||Q^{B})+(\alpha-1)D_{\alpha}(P||Q^{B})$ (49)
+$=(1-\alpha)\left[D(P||Q^{B})-D_{\alpha}(P||Q^{B})\right].$ (50)
 Finally, we substitute this result back into the original expression for expected work (3):
-$\mathbb{E})[W]=D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A,r}X)\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})\\ =\alpha D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})$ (54)
-Thus, the expected work extracted for the optimal strategy corresponding to risk parameter $r$ is:
-$\boxed{\mathbb{E}[W]=\alpha D(P_{X}||Q^{B}_{X})k_{B}T+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})k_{B}T,\quad\text{where}\quad\alpha=\frac{1}{1+r}}$ (55)
-Appendix G Conditions for rationality for r<1r<-1
-In section eq. 41 we showed that the certainty equivalent for work extraction for a rational agent with CARA utility function is a Rényi divergence:
-$W_{CE}=D_{\alpha}(P_{X}||Q_{X}^{B})k_{B}T\quad\text{where}\quad\alpha=\frac{1}{1+r}$ (56)
-The work extracted in a specific outcome $x$ is:
-$W(x)=\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)k_{B}T$
-where $Q_{X}^{A,r}$ is the optimal strategy for a given value of the risk parameter $r$.
-The violation of rationality (First-Order Stochastic Dominance) occurs if the certainty equivalent is more negative than the worst-case work output:
-$W_{CE}<\min_{x}W(x)$
-Substituting in the expression for the certainty equivalent, we get:
-$D_{\alpha}(P_{X}||Q_{X}^{B})<\min_{x}\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)$
-This is the general condition. We can now use the specific form of $Q_{X}^{A,r}$ to simplify the right-hand side. Recall:
-$Q_{X}^{A,r}(x)=\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}}$
-Defining $Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}$ as the normalisation constant and substituting into the work expression:
-$W(x)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{Z(\alpha)Q_{X}^{B}(x)}\right)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{-\alpha}}{Z(\alpha)}\right)=\alpha\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
-To find the minimum value of this expression, note that the term $\ln Z(\alpha)$ is independent of $x$. Therefore:
-$\min_{x}W(x)=\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
-The normalisation constant $\ln Z(\alpha)$ is directly related to the Rényi divergence by:
-$D_{\alpha}(P_{X}||Q_{X}^{B})=\frac{1}{\alpha-1}\ln Z(\alpha)$
-Now, we can substitute back into our violation condition (Eq. 1):
-$D_{\alpha}(P_{X}||Q_{X}^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
-Substitute $D_{\alpha}$ and rearrange:
-$\frac{1}{\alpha-1}\ln Z(\alpha)<\alpha\cdot M-\ln Z(\alpha)\quad\text{where}\quad M=\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$
-$\ln Z(\alpha)\left(\frac{1}{\alpha-1}+1\right)<\alpha M$
-$\ln Z(\alpha)\left(\frac{1+(\alpha-1)}{\alpha-1}\right)<\alpha M$
-$\frac{\ln Z(\alpha)\cdot\alpha}{\alpha-1}<\alpha M$
-For $\alpha<0$ (which is true for r<1r<-1, the case we are interested in), dividing both sides by $\alpha$ flips the sign of the inequality:
-$\frac{\ln Z(\alpha)}{\alpha-1}>M$
-Since the left-hand side is exactly the Rényi divergence $D_{\alpha}(P_{X}||Q_{X}^{B})$, the condition for a violation of rationality simplifies to:
-$\boxed{D_{\alpha}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)}$
-since this never holds, rationality is never violated.
+$D(P||Q^{B})-D(P||Q^{A,r})$ $=D(P||Q^{B})-(1-\alpha)\left[D(P||Q^{B})-D_{\alpha}(P||Q^{B})\right]$ (51)
+$=D(P||Q^{B})-(1-\alpha)D(P||Q^{B})+(1-\alpha)D_{\alpha}(P||Q^{B})$ (52)
+$=\alpha D(P||Q^{B})+(1-\alpha)D_{\alpha}(P||Q^{B}).$ (53)
+Thus, the expected work extracted for the optimal strategy corresponding to risk parameter $r$ is given by Eq. (9).
+Appendix F Conditions for rationality for r<1r<-1
+In Eq. (11), we showed that the certainty equivalent for work extraction for a rational agent with CARA utility function is a Rényi divergence. The work extracted in a specific outcome $x$ is
+$w(x)=k_{B}T\ln\left(\frac{Q^{A,r}(x)}{Q^{B}(x)}\right),$ (54)
+where $Q^{A,r}$ is the optimal strategy for a given value of the risk parameter $r$. Substituting this to the certainty equivalent, we get
+$D_{\alpha}(P||Q^{B})<\min_{x}\ln\left(\frac{Q^{A,r}(x)}{Q^{B}(x)}\right).$ (55)
+This is the general condition. We can now use the specific form of $Q^{A,r}$ to simplify the right-hand side. From Eq. (42) and Eq. (48), we obtain
+$w(x)=\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)Q^{B}(x)}\right)=\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{-\alpha}}{Z(\alpha)}\right)=\alpha\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}).$ (56)
+To find the minimum value of this expression, recall that $\alpha<0$ for r<1r<-1. Therefore,
+$\min_{x}w(x)=\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}).$ (57)
+We substitute this back into our violation condition (55) and obtain
+$D_{\alpha}(P||Q^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}),$ (58)
+or equivalently,
+$\alpha D_{\alpha}(P||Q^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right).$ (59)
+Again noticing that $\alpha<0$, the violation condition (55) reduces to
+$D_{\alpha}(P||Q^{B})>\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right),$ (60)
+which is Eq. (15). Since this never holds, rationality is never violated.
+Appendix G Derivation of the work bound
+In this Appendix, we show the derivation of Eq.
+(23).
+Recall that the optimal distribution $Q^{A*,\mu}(x)$ is given by Eq.
+(22), for which we define $Z(\mu)=\sum_{x^{\prime}}P(x^{\prime})^{\mu}Q^{B}(x^{\prime})^{1-\mu}$.
+We can show that
+$D(Q^{A*,\mu}\|Q^{B})=\mu\sum_{x}Q^{A*,\mu}\ln\frac{P(x)}{Q^{B}(x)}-\ln Z(\mu),$ (61)
+$D(Q^{A*,\mu}\|P)=-(1-\mu)\sum_{x}Q^{A*,\mu}\ln\frac{P(x)}{Q^{B}(x)}-\ln Z(\mu).$ (62)
+By noting that $D_{\mu}(P\|Q^{B})=-\frac{1}{1-\mu}\ln Z(\mu)$, we obtain
+$D(Q^{A*,\mu}\|Q^{B})=D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}D(Q^{A*,\mu}\|P).$ (63)
+Now recall that $Q^{A*,\mu}=\lambda_{x^{n}}$. Because of the risk constraint (21),
+$D(Q^{A*,\mu}\|Q^{B})\geq D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}\frac{1}{n}\ln\frac{1}{\epsilon}.$ (64)
+Going back to the work expression (18), we obtain
+$W_{n}(x^{n})\geq k_{B}TnD(Q^{A*,\mu}||Q^{B})\geq k_{B}Tn\left(D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}\frac{1}{n}\ln\frac{1}{\epsilon}\right),$ (65)
+which proves Eq. (23).